### v1

# Neurons Speak in Ranges: Breaking Free from Discrete Neuronal Attribution

Abstract
Interpreting and controlling the internal mechanisms of large language models (LLMs) is crucial for improving their trustworthiness and utility. Recent efforts have primarily focused on identifying and manipulating neurons by establishing discrete mappings between neurons and semantic concepts. However, such mappings struggle to handle the inherent polysemanticity in LLMs, where individual neurons encode multiple, distinct concepts. This makes precise control challenging and complicates downstream interventions. Through an in-depth analysis of both encoder and decoder-based LLMs across multiple text classification datasets, we uncover that while individual neurons encode multiple concepts, their activation magnitudes vary across concepts in distinct, Gaussian-like patterns. Building on this insight, we introduce $\mathsf{NeuronLens}$, a novel range-based interpretation and manipulation framework that provides a finer view of neuron activation distributions to localize concept attribution within a neuron. Extensive empirical evaluations demonstrate that $\mathsf{NeuronLens}$ significantly reduces unintended interference, while maintaining precise control for manipulation of targeted concepts, outperforming existing methods.
Polysemanticity,Interpretability
1 Introduction
Large language models (LLMs) have shown unprecedented performance in various tasks involving natural language understanding, generation, and transformation (Brown et al., 2020; Bommasani et al., 2022; Touvron et al., 2023; Raffel et al., 2019). However, the inner workings of LLMs remain largely opaque (Burkart & Huber, 2021), as their representations are distributed across billions of parameters. This lack of interpretability raises critical concerns about reliability, fairness, and trustworthiness, particularly in high-stakes domains such as healthcare, law, and education. To this end, neuron-level interpretability can address these concerns by enabling researchers to uncover how individual neurons encode semantic concepts and contribute to model outputs. With this understanding, researchers can diagnose safety risks (Wei et al., 2024; He et al., 2024), refine model outputs (Meng et al., 2023; Rizwan et al., 2024), optimize efficiency through pruning (Frankle & Carbin, 2019; Haider & Taj, 2021), and steer models toward desired objectives (Subramani et al., 2022).
Recent research efforts have made significant progress in neuron-level interpretation by identifying salient neurons that influence model behavior (Dalvi et al., 2019a; Antverg & Belinkov, 2022; Conmy et al., 2023; Marks et al., 2024). Approaches such as maximal activation analysis identify neurons that exhibit the highest activation values for specific concepts (Foote et al., 2023; Frankle & Carbin, 2019). Probe-based methods employ auxiliary classifiers to distinguish between concepts using hidden representations (Dalvi et al., 2019a, b), while probeless approaches bypass the need for classifiers by directly analyzing neurons, improving simplicity and interpretability (Antverg & Belinkov, 2022). Other techniques include circuit discovery, which attributes concepts to groups of interacting neurons (Olah et al., 2020; Conmy et al., 2023), and causal analysis, which intervenes on internal components to identify their role in model behavior (Vig et al., 2020).
Figure 1: $\mathsf{NeuronLens}$ leverages distinct, Gaussian-like activation patterns of neurons to enable fine-grained concept attribution and selective concept manipulation with minimal interference with other concepts.
While these methods have advanced our understanding, they often rely on discrete neuron-to-concept mappings, which assume that entire neurons encode single concepts. However, neurons frequently exhibit polysemanticity – the ability to encode multiple, seemingly unrelated concepts (Lecomte et al., 2024; Marshall & Kirchner, 2024). Given this heterogeneous encoding of concepts, traditional approaches often lead to unintended consequences when manipulating neurons, as changes intended for one concept may inadvertently affect others encoded by the same neuron or suboptimal interpretations of concepts (Sajjad et al., 2022).
Despite being traditionally viewed as a challenge, could polysemanticity instead provide a unique lens for advancing interpretability and model control? If individual neurons encode multiple concepts, might their activation spectrum reveal distinct and identifiable patterns for each concept? Could these patterns enable precise interventions that adjust one concept while minimizing interference with others, overcoming the limitations of coarse, monolithic neuron-to-concept mappings?
This work seeks to address these questions by analyzing the activation patterns of neurons in both encoder-based and decoder-based LLMs. Through statistical and qualitative analysis, we find that neuronal activations for concepts follow Gaussian-like distributions, with distinct patterns for different concepts. Our key insight is that the true fundamental unit of interpretability lies at a level more fine-grained than the neuron itself. Within a neuron’s activation spectrum, activation ranges corresponding to specific concepts emerge as the fundamental unit. This novel perspective enables a more precise approach to neuron interpretation and manipulation, addressing the limitations of traditional, discrete neuron-to-concept mappings.
Building upon these insights, we introduce $\mathsf{NeuronLens}$, a novel range-based framework for neuron interpretation and manipulation. $\mathsf{NeuronLens}$ maps specific activation ranges within a neuron’s distribution to individual concepts, rather than attributing entire neurons to single concepts. For each concept, $\mathsf{NeuronLens}$calculates a range that covers approximately 98.76% of the activation spectrum, capturing the concept-specific activations while excluding unrelated concepts. Through experiments on encoder-based and decoder-based LLMs across several text classification datasets, we show that $\mathsf{NeuronLens}$. significantly reduces unintended interference by up to 25 percentage points, while maintaining precise control for manipulation of targeted concepts, outperforming existing methods.
In summary, this work makes the following contributions:
  • We uncover that neuronal activations in LLMs form Gaussian-like distributions, with salient neurons exhibiting limited overlap in their activation patterns across concepts.
  • We show that activation ranges within a neuron’s activation spectrum are the true fundamental unit of interpretability, offering a precise framework for neuron-level analysis.
  • We propose $\mathsf{NeuronLens}$, a range-based framework for interpreting and manipulating neuronal activations, which enables fine-grained concept attribution and reduces unintended interference.
2 Neuron Interpretation Analysis
2.1 Preliminaries
Neuron. We refer to the output of an activation as a neuron. In a transformer model, we consider neurons of hidden state vectors of different transformer layers. Formally, given a hidden state vector $\mathbf{h}^{l}\in\mathbb{R}^{d}$ of size $d$ produced by layer $l$, $h^{l}_{j}$ denotes its $j$-th neuron, i.e., the $j$-th component of $\mathbf{h}^{l}$.
Concept. A concept $c\in C$ is a high-level semantic category that groups each input instance (or components of every instance), where $C$ is the set of all concepts. For example, in a language task, a sentence can be categorized into 4 types: declarative, interrogative, imperative, and exclamatory, where each type is a concept. Words of a sentence can also have concepts like noun, verb, adjective, adverb, etc. In this study, we focus on the situation that all input samples are labeled with concepts.
Saliency Ranking. A saliency ranking orders the importance of neurons based on some saliency metric. For a hidden state vector $\mathbf{h}^{l}\in\mathbb{R}^{d}$, $s_{j,c}$ denotes the value of the saliency metric for the $j$-th neuron with respect to a concept $c$. The saliency ranking $(r_{c}(1),r_{c}(2),\cdots,r_{c}(d))$ is a permutation of the indices of neurons $(1,2,\cdots,d)$, where $r_{c}(j)<r_{c}(i)$ if $s_{j,c}>s_{i,c}$. The saliency metric is usually predetermined, e.g., the absolute value of each neuron.
Concept Learning. Given a hidden state vector $\mathbf{h}^{l}$ as input, the associated concept can be the output of an appended neural network (e.g., several fully connected layers). The parameters of this appended neural network can be trained using training samples labeled with concepts.
Table 1: Performance drops relative to Baseline configuration(i.e: unaltered model’s performance) for three techniques: Probeless, Probe, and Max. All values show the difference from Base values. Results are for Emotions dataset on the GPT-2 model. 30% neurons used. The trimmed mean excludes the top and bottom 10% of values.
Probeless Probe Max
Class Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
Class 0 -0.738 -0.733 -0.103 -0.097 -0.613 -0.650 -0.010 -0.038 -0.695 -0.751 -0.125 -0.124
Class 1 0.045 0.041 -0.113 -0.112 -0.014 -0.015 -0.010 -0.034 -0.879 -0.882 -0.019 -0.009
Class 2 -0.570 -0.541 -0.052 -0.057 0.017 0.009 -0.347 -0.359 -0.776 -0.736 -0.029 -0.032
Class 3 -0.164 -0.166 -0.035 -0.038 0.078 0.061 -0.047 -0.104 -0.713 -0.714 -0.006 -0.007
Class 4 -0.623 -0.617 -0.087 -0.084 -0.005 -0.010 -0.003 -0.020 -0.754 -0.753 -0.240 -0.248
Class 5 -0.817 -0.714 -0.101 -0.105 -0.206 -0.127 0.003 -0.010 -0.587 -0.601 -0.301 -0.308
2.2 Concept Erasure
To assess the performance of neuronal ranking obtained using attribution methods, concept erasure acts as a critical diagnostic intervention to determine the causal effect of identified neurons for a given concept (Dalvi et al., 2019b; Dai et al., 2022; Dalvi et al., 2019c; Morcos et al., 2018). The core idea is that by manipulating the salient neurons, it is possible to precisely eliminate a specific identified concept while causing minimal disruption to the other concepts learned by the model. This can be formalized as follows: given a concept-learning model $M$ that maps any input instance $x$ (or part of an instance) to a concept $M(x)\in C$, an ideal intervened model $M^{\prime}_{\text{ideal}}$ after erasing a target concept $c\in C$ should satisfy the following property:
$M^{\prime}_{\text{ideal}}(x)\begin{cases}\neq M(x)\quad\text{ if}M(x)=c,\\ =M(x)\quad\text{ if }M(x)\neq c.\end{cases}$
A popular approach of concept erasure in neuronal analysis literature (Dai et al., 2022; Antverg & Belinkov, 2022) is zeroing out specific neurons that are “important” to the target concept.
2.3 Salient Neurons Extraction
Problem setup and preparation: We use a transformer model to process sentences. Before doing neuron interpretation, we record activations for training samples of different concepts as preparation. Specifically, if we want to interpret neurons of $\mathbf{h}^{l}$ (hidden state vector at layer $l$), we will traverse the training dataset and store the values of $\mathbf{h}^{l}$ and the associated concepts of all samples into a set $H^{l}$. The set $H^{l}$ is further partitioned into $H^{l}_{c}$ for all concepts $c\in C$. Such preparation is commonly used in the literature on neuron interpretation (Dalvi et al., 2019c, b; Antverg & Belinkov, 2022).
Max activations. Frankle & Carbin (2019) extract high neural activations as a saliency ranking metric relying upon the rationale that maximally activating neurons are salient as these neurons play a critical role in controlling the model’s output, highlighting their importance for a concept $c$. Probe analysis. Dalvi et al. (2019b) train a linear classifier on the hidden representations $H^{l}_{c}$ to distinguish between concepts. The learned model weights are then utilized as a saliency ranking. Probeless. Antverg & Belinkov (2022) examine individual neurons, without the need for auxiliary classifiers, using the element-wise difference between mean vectors. Details of these approaches are provided in Appendix C.
Mutual comparison of these approaches is provided in Table 1. From the table, we observed that irrespective of the method used to obtain saliency ranking, a single concept eraser using salient neurons causes deterioration in performance across several concepts. Max activation causes the highest degree of deterioration in the targeted concept while maintaining a comparable deterioration in auxiliary concepts, based on this finding, we adopt max activation ranking for saliency ranking. Moreover, we hypothesize that one reason for such deterioration in overall performance is due to the polysemantic nature of neurons.
Figure 2: Illustrates the percentage overlap of salient neurons identified through maximal activation ranking across concepts (i.e., classes), in all combinations, for the GPT-2 model trained on different datasets. Percentage of salient neurons extracted is 30%.
3 Polysemanticity
The polysemanticity of neuronal units, salient neurons encode information about multiple concepts, pose a challenge to neural network interpretation and manipulation. In this section, we discuss the degree of polysemanticity in salient neurons in detail.
Figure 3: Neuronal Activation Patterns of six neurons on AG-News dataset. Two on the top are the highest activating neurons, two in the middle are middle-ranked neurons, and two at the bottom are the lowest activating neurons for class 1.
Superposition. Polysemanticity (i.e., superposition of neurons) is primarily thought to arise when a model must represent significantly more features than the available capacity of its representational space or due to the training paradigms used. Limited representation space forces to encode multiple unrelated features within the same neuron to achieve high performance (Anthropic, 2023). Additionally, training paradigms such as subtokenization in tokenization schemes, which aim to reduce vocabulary size to decrease the model size and improve generalization for tokens with the low natural frequency. This results in context-dependent token decomposition, where a single token may be split differently based on its surrounding context (Sennrich et al., 2016). This contextual nature leads to a state of superposition, where multiple semantic meanings are encoded within the same neural activations (Elhage et al., 2022; Meng et al., 2022). Weight initialization can also contribute to polysemanticity, as demonstrated empirically by (Lecomte et al., 2024). Their findings reveal that neurons exhibiting activation patterns equidistant from or similarly close to multiple conceptual representations at initialization can lead to polysemanticity, even when the network’s capacity (i.e., the number of neurons) exceeds the number of concepts to be represented.
Figure 4: Neuronal Activation Patterns comparison of neurons 480 and 675. The plots show class-specific activity patterns with fitted Gaussian curves. Both neurons were part of salient set of all classes when top 5% salient selected on AG-News dataset.
Polysemanticity in salient neurons. Given that salient neurons have a strong causal association with the concept of interest, their tendency to be mono-semantic should be high, but we find that there is a high degree of polysemanticity in salient neurons.
We investigate this by extracting 30% salient neurons (i.e.: Max neurons) for different datasets on the GPT-2 model. The results show that there is a considerable overlap of salient neurons between concepts (classes) as shown in Figure 2. In the case of a two-class dataset IMDB, the overlap of salient neurons, selected by max, is more than 60%. This shows a high degree of polysemanticity.
Consequently, we extrapolate that salient neural representations may exist in a superimposed configuration, wherein a subset of the salient neurons encode information through intricate, overlapping activation patterns. The monolithic attribution paradigm potentially oversimplifies the complex, distributed nature of neuronal activation as can be seen in polysemanticity (Lecomte et al., 2024; Marshall & Kirchner, 2024) where single neuron learns multiple seemingly unrelated concepts and elucidates them at different activation values
4 Neuronal Activation Patterns
In this section, we analyze the properties of neuronal activations of the salient neurons (including polysemantic) extracted via maximal activation. Our findings indicate that neuronal activations on a concept form a Gaussian-like distribution. We also find that salient neurons have a distinct Gaussian distribution of activations for different concepts with minimal overlap with other concept activations.
Qualitative Evaluation.
Figure 5: Box plot of neural activation of 11 polysemantic neurons (i.e: neurons in the salient group for all classes, percentage selected: 5% top salient) for 4 randomly selected classes out of 14 classes of DBPedia-14 dataset.
To visually demonstrate that neuron activations for a concept $c$ follow a Gaussian-like distribution, we extract model representations as described in Section 2.1. Using saliency ranking $r_{c}$ for a single concept class, we examine neurons from different ranking positions in the GPT-2 model on the AG-News dataset: two top-ranked neurons ($r_{c}\leq 2$), two middle-ranked neurons ($r_{c}\approx d/2$), and two bottom-ranked neurons ($r_{c}\geq d-1$). In Figure 3 and 4 we use Kernel Density Estimation (KDE) to visualize these distributions. Figure 3 reveals that while the activations are Gaussian-like for different concepts, salient neurons demonstrate distinct activation patterns with limited overlap, middle-ranked neurons show a higher degree of overlap than the top ones, whereas non-salient neurons (bottom two) exhibit the highest overlap in their activation distributions. On GPT-2 model.
Additionally in Figure 4, we identify and visualize two distinct types of polysemantic neurons that appear in the salient sets across all classes, when 5% salient set was selected, in the dataset. The first type, exemplified by neuron 480, maintains partially separable activation patterns despite being polysemantic, suggesting some degree of class-specific behavior. In contrast, the second type, represented by neuron 675, exhibits completely overlapping activation patterns across all classes, making it hard to disentangle. To further investigate this phenomenon, Figure 5 presents a broader analysis of neurons from the polysemantic subset, identified using a 5% saliency threshold. By examining these neurons’ behavior across four randomly selected classes (out of 14 total classes), we observe that most polysemantic neurons exhibit a high degree of separability, for some classes – while they respond to multiple classes, they tend to operate in partially separable activation ranges, supporting the possibility of meaningful disentanglement.
Table 2: Statistical analysis of different datasets showing skewness, kurtosis, and Kolmogorov-Smirnov test results.
Dataset Skewness Kurtosis KS-Test
stanfordnlp/imdb 0.0014 3.6639 1.0000
fancyzhx/dbpedia_14 -0.0007 3.9360 0.9782
dair-ai/emotion 0.0015 3.0198 0.9446
fancyzhx/ag_news -0.0013 3.2060 0.9918
stanfordnlp/sst2 -0.0083 3.2038 1.0000
Quantitative Evaluation. To quantify the effect of Gaussian-like distribution of neurons, for $c\in C$, we perform statistical analysis of activations. We computed the skewness, kurtosis (Joanes & Gill, 1998) and analyzed the normality of neuronal activations using Kolmogorov-Smirnov (KS) test (Massey Jr, 1951). The results for basic distributional properties across all neurons are presented in Table 2. The average skewness is close to 0 across all datasets, indicating strong symmetry (ideal normal distribution: 0), and the average kurtosis is close to 3, nearly identical to the expected value for a normal distribution (3.0), on GPT-2 model.
To quantitatively assess normality, while accounting for practical significance, we employ the KS test with an effect size threshold of 10%. This approach tests whether the distribution remains within a reasonable bound of normality, rather than testing for perfect normality, which is overly strict for real-world data. For each neuron, we normalize the activations to zero mean and unit variance, then compute the KS statistic against a standard normal distribution. The KS statistic represents the maximum absolute difference between the empirical and theoretical cumulative distribution functions. Using a threshold of 0.1 (allowing maximum 10% deviation from normal), we find that close to 100% of the neurons exhibit practically normal distributions. The combination of near-ideal skewness and kurtosis values, visual confirmation through KDEs, and our effect size-based KS tests provide strong evidence that the activations follow approximately normal distributions.
Table 3: Evaluation of selected models on IMDB, SST2, AG-News, and DBPedia-14 datasets using activation range and neuron masking techniques. Performance metrics, calculated using trimmed means at the class level, include class accuracy (Acc), class prediction probability (Conf), average accuracy across other classes (CAcc), and average class prediction probability across other classes (CConf). Base Values represent baseline model performance, Activation Range Masking and Neuron Masking entries show the drop from these baselines. The trimmed mean is calculated by removing extreme values (top and bottom 10%) before computing the average to reduce the impact of outliers. We only bold the drops in auxiliary concepts (CAcc, CConf). The Drops in targeted concepts are comparable with Neuron Masking with slightly better results overall. For Bert and GPT-2 50% and for Llama-3.2-3B 30% neurons selected.
Model Dataset Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
BERT IMDB 0.928 0.904 0.928 0.904 -0.190 -0.353 0.059 -0.078 -0.184 -0.360 0.058 0.030
SST2 0.910 0.903 0.910 0.903 -0.051 -0.313 0.031 -0.046 -0.060 -0.330 0.031 0.043
AG-NEWS 0.948 0.929 0.948 0.929 -0.271 -0.590 0.012 -0.074 -0.261 -0.590 0.013 -0.009
Emotions 0.894 0.834 0.917 0.876 -0.291 -0.633 0.013 -0.265 -0.279 -0.635 0.014 -0.069
DBPedia-14 0.992 0.991 0.990 0.989 -0.028 -0.786 $0$0.017 -0.015 -0.766 $0$0.000
GPT-2 IMDB 0.952 0.939 0.952 0.939 -0.196 -0.188 0.033 0.045 -0.195 -0.197 0.031 0.042
SST2 0.966 0.958 0.966 0.958 -0.165 -0.190 0.025 0.032 -0.159 -0.192 0.025 0.028
AG-NEWS 0.945 0.933 0.945 0.933 -0.871 -0.877 -0.155 -0.163 -0.849 -0.862 -0.063 -0.223
Emotions 0.905 0.892 0.930 0.919 -0.735 -0.738 -0.103 -0.103 -0.737 -0.739 -0.044 -0.046
DBPedia-14 0.993 0.990 0.990 0.988 -0.810 -0.845 -0.154 -0.177 -0.782 -0.825 -0.015 -0.031
Llama-3.2-3B IMDB 0.952 0.939 0.952 0.939 -0.196 -0.188 0.033 0.045 -0.195 -0.197 0.031 0.042
SST2 1.000 0.559 1.000 0.559 -0.760 -0.429 -0.394 -0.295 -0.756 -0.427 -0.384 -0.291
AG-NEWS 1.000 0.744 1.000 0.744 -0.934 -0.725 -0.660 -0.572 -0.935 -0.725 -0.484 -0.454
Emotions 0.815 0.472 0.823 0.477 -0.795 -0.470 -0.696 -0.429 -0.797 -0.469 -0.594 -0.404
DBPedia-14 1.000 0.533 1.000 0.563 -0.992 -0.528 -0.912 -0.445 -0.986 -0.527 -0.663 -0.354
5 Activation Ranges-guided Intervention
Given that the neuronal activations are Gaussian-like with separable means, we can interpret and manipulate the neurons in a more precise manner instead of zeroing out the neurons. Specifically, we only zero out a salient neuron (by saliency ranking) when its value is within a certain range. The core part of this design is selecting a suitable range that is highly correlated to the target concept $c$ that we want to remove. Intuitively, this range-based method conducts fine grained manipulation of neurons and thus alleviates the negative impact on other non-target concepts.
In this work, we adopt a straightforward choice of the aforementioned range, which is easy to calculate and implement. Specifically, we first calculate the empirical average $\mu\in\mathds{R}$ and standard deviation $\sigma\geq 0$ of the values of the salient neuron for all samples associated with the target concept $c\in C$. After that, we assign the range as $[\mu-\tau\times\sigma,\ \mu+\tau\times\sigma]$, where $\tau>0$ is a hyperparameter to make a tradeoff between erasing the target concept $c$ (using larger $\tau$) and smaller impact on other concepts (using smaller $\tau$). For this work, we use $\tau=2.5$. We chose this value to cover the majority of the target concept activations to produce a similar deterioration of the targeted concept.
Mathematically, our range-based method can be described as follows. To erase a target concept $c$ for a sample $x$ (e.g., a sentence or a word), we manipulate the salient neuron $h_{j}^{l}(x)$ ($j$-th element of the hidden state vector $\mathbf{h}^{l}(x)\in\mathds{R}^{d}$ at layer $l$).
$h_{j}^{l}(x)\begin{cases}=0&\text{ if }h_{j}^{l}(x)\in\text{CorrelatedRange}(l,j,c),\\ =h_{j}^{l}(x)&\text{ otherwise},\end{cases}$
where in this work we let
$\text{CorrelatedRange}(l,j,c)=[\mu-2.5\sigma,\ \mu+2.5\sigma],$
$\mu=\frac{1}{\left|H^{l}_{c}\right|}\sum_{\mathbf{h}^{l}\in H^{l}_{c}}h_{j}^{l},\ \quad\sigma=\sqrt{\frac{1}{\left|H^{l}_{c}\right|}\sum_{\mathbf{h}^{l}\in H^{l}_{c}}\left(h_{j}^{l}-\mu\right)^{2}}.$
Notice that $H^{l}_{c}$ was defined in problem setup and preparation of Section 2.3, which denotes the set of hidden state vector $\mathbf{h}^{l}(x_{c})$ at layer $l$ for all training samples $x_{c}$ associated with concept $c$. Here ||$\cdot$|| ⋅ | denotes the cardinality of a set.
Table 4: Evaluation of selected models on the AG-News dataset using neuron and range masking techniques. Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance. For Bert and GPT-2 50% and for Llama-3.2-3B 30% neurons selected.
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
BERT Class 0 0.945 0.936 0.949 0.927 -0.205 -0.587 0.004 -0.076 -0.198 -0.589 0.007 -0.010
Class 1 0.993 0.988 0.933 0.910 -0.225 -0.659 0.004 -0.077 -0.194 -0.650 0.003 -0.012
Class 2 0.905 0.881 0.962 0.945 -0.300 -0.536 0.014 -0.079 -0.298 -0.542 0.014 -0.009
Class 3 0.949 0.913 0.948 0.935 -0.354 -0.577 0.026 -0.065 -0.353 -0.579 0.025 -0.005
GPT-2 Class 0 0.955 0.951 0.941 0.928 -0.920 -0.926 -0.231 -0.224 -0.919 -0.925 -0.019 -0.008
Class 1 0.986 0.981 0.931 0.917 -0.926 -0.931 -0.253 -0.257 -0.912 -0.916 -0.054 -0.069
Class 2 0.897 0.886 0.960 0.949 -0.696 -0.737 -0.110 -0.132 -0.678 -0.725 -0.097 -0.306
Class 3 0.940 0.916 0.946 0.939 -0.940 -0.916 -0.024 -0.037 -0.887 -0.882 -0.080 -0.510
Llama-3.2-3B Class 0 1.000 0.936 1.000 0.680 -0.995 -0.934 -0.530 -0.427 -0.995 -0.934 -0.345 -0.306
Class 1 1.000 0.742 1.000 0.744 -0.870 -0.680 -0.615 -0.599 -0.875 -0.681 -0.515 -0.503
Class 2 1.000 0.655 1.000 0.773 -0.895 -0.646 -0.795 -0.634 -0.895 -0.646 -0.655 -0.549
Class 3 1.000 0.642 1.000 0.778 -0.975 -0.641 -0.698 -0.630 -0.975 -0.640 -0.420 -0.459
5.1 Experimental Setup
Models. This study employs a range of transformer-based architectures, including encoder and decoder models such as fine-tuned BERT (Devlin et al., 2019), DistilBERT (Sanh et al., 2020), GPT-2 (Radford et al., 2019), and pretrained Llama-3.2-3B (Grattafiori, 2024). We incorporate our methodology at the penultimate layer, abiliation for layer selection is provided in the Appendix E. The training details for the models are provided in appendix B.
Datasets. We consider various classification based tasks; sentiment analysis (IMDB, Maas et al. (2011)), (SST2, Socher et al. (2013)), emotion detection (Dair-Ai/Emotions Saravia et al. (2018)), news classification (AG-News (Zhang et al., 2015)) and article content classification tasks (DBPedia-14 (Zhang et al., 2015)).
Metrics. For performance evaluation, we employ two quantitative metrics: prediction accuracy and the model’s confidence score for the ground truth. Initially, we established baseline measurements of both accuracy and confidence for all concepts $C$ in the unmodified model. Post-intervention measurements are recorded for both the target concept $c$ and auxiliary concepts $c^{\prime}$. The effectiveness and precision of neuron identification can be assessed through two key metrics: (1) the magnitude of performance degradation for concept $c$, and (2) the extent of unintended impact on auxiliary concepts $c^{\prime}$. Throughout our analysis, we denote the accuracy and confidence metrics for concept $c$ as Acc and Conf respectively, while corresponding measurements for auxiliary concepts $c^{\prime}$ are represented as CAcc and CConf.
5.2 Results and Analysis
Table 3 presents comprehensive results across five datasets, demonstrating the effectiveness of our range-based masking approach compared to traditional neuron masking. Range-based masking better preserves auxiliary concepts across all settings performance (CAcc, CConf), suggesting more precise concept isolation.
Multi-class classification tasks (AG-News, Emotions, DBPedia-14) demonstrate more pronounced intervention effects. Range masking achieves substantial primary task degradation while maintaining auxiliary concept performance, particularly evident in AG-News results (Table 4). The class-level analysis reveals asymmetric impacts: Classes 0 and 1 show larger performance drops (-0.919, -0.912 for GPT-2) compared to Classes 2 and 3 (-0.678, -0.887), suggesting varying degrees of concept distribution across neurons.
Model architecture and training methodology significantly influence intervention outcomes. BERT, trained directly on the classification tasks with a dedicated classification head, demonstrates remarkable resilience with minimal accuracy degradation (-0.015 on DBPedia-14) overall, under range masking it maintains auxiliary concept performance. This robustness likely stems from its task-specific fine-tuning, which enables more distributed and redundant concept representations. In contrast, GPT-2, despite being fine-tuned but trained in an autoregressive manner, shows substantially higher vulnerability with major drops in AG-NEWS ($\Delta_{acc}=-0.849$) and DBPedia-14 ($\Delta_{acc}=-0.782$). This increased sensitivity may be attributed to its autoregressive training objective, which potentially leads to more sequential and less redundant concept encodings. Llama-3.2-3B, evaluated in a few-shot setting without task-specific training, experiences the most severe degradation across all datasets (often exceeding -0.900), suggesting that pre-trained representations without task-specific fine-tuning are more vulnerable to targeted neuron interventions. These patterns indicate that task-specific training, particularly with classification-oriented objectives, may foster more robust and distributed concept representations that are more resilient to neuron interventions.
The effectiveness of range-based masking in preserving auxiliary concepts while achieving targeted performance degradation suggests that concept encoding in neural networks may follow activation range patterns rather than being strictly neuron-specific. This is particularly evident in the AG-News results, where range masking maintains better auxiliary concept scores (CAcc, CConf) compared to neuron masking across all classes and models.
Figure 6: Accuracy comparison between Neuronal Range manipulation (green) and complete neuron manipulation (orange) methods. Both approaches show similar primary task degradation (solid lines) but differ significantly in auxiliary task stability (dashed lines). The Neuronal Range method maintains more stable auxiliary task performance than Maximal Activation’s steep degradation after 50%. Model GPT-2
On binary classification tasks (IMDB, SST2), both masking approaches show moderate performance drops in targeted concepts, with BERT experiencing accuracy decreases of -0.184 and -0.060 respectively under range masking. The relatively modest impact on primary task performance, coupled with minimal auxiliary concept interference, indicates significant neuronal redundancy in binary classification tasks multiple neurons appear to encode similar sentiment-related features.
Percentage Masking Effect As the proportion of masked neurons increases, the performance improvement over the maximum activation baseline becomes more pronounced. Beyond a critical masking threshold, the baseline performance deteriorates exponentially, while our proposed method maintains stable performance up to 100% masking. This phenomenon can be explained by two key factors. First, as we ablate a higher percentage of neurons, the probability of encountering polysemantic neurons increases as a statistical consequence. Second, neurons identified through maximum activation are more likely to exhibit discriminatory behavior between concepts or demonstrate strong conceptual associations. when analyzing neurons with lower activation patterns (not strongly associated with concepts), discrete classification becomes an increasingly crude metric, as evident from Figure 6. After 50% masking, we see a steep drop in the general performance of the model, in the case of neuron masking. This suggests the need for a more nuanced approach to mapping information as to neural activations as blocking/manipulating a higher percentage of the model’s representation creates a significant deviation from the original model’s behavior. Our range-based methodology provides such precision, enabling more accurate characterization of neuronal behavior across the activation spectrum.
6 Related Work
While we have discussed closely related approaches in Section 2, here, we briefly review additional relevant techniques. Circuit discovery identifies groups of neurons that jointly encode concepts, providing a structured view of model behavior (Marks et al., 2024; Conmy et al., 2023; Olah et al., 2020). However, extracting circuits is computationally intensive and lacks fine-grained neuron-level attribution. Gradient-based methods attribute predictions to input features by tracking gradients through the network, with integrated gradients (Sundararajan et al., 2017; Dai et al., 2022) being a widely used approach. However, they struggle with polysemanticity, as they do not disentangle overlapping concepts within neurons. Causal analysis methods intervene on internal components to assess their role in encoding concepts. Causal tracing measures the effect of corrupting activations on model performance (Vig et al., 2020; Meng et al., 2022), while causal mediation analysis quantifies information propagation through neurons (Vig et al., 2020). Although effective, these techniques require costly perturbation experiments. Beyond neuron-level analysis, representation-level methods examine hidden states and their relationship to model outputs. Logit lens (nostalgebraist, 2020) and tuned lens (Belrose et al., 2023) project intermediate representations into the model’s vocabulary space to track information flow, while sparse probing (Gurnee et al., 2023) compresses hidden representations into sparse, interpretable subspaces. While prior work has advanced interpretability, most methods rely on discrete neuron-to-concept mappings, which fail to account for polysemanticity. Our work extends activation-based approaches by introducing activation ranges as the fundamental unit of interpretability to enable more precise concept attribution and intervention.
7 Conclusion
In this work, we challenged traditional assumptions about neuron interpretability by reframing polysemanticity as a resource rather than a limitation in interpreting neurons. Through an in-depth analysis of encoder- and decoder-based large language models across multiple text classification datasets, we uncovered that neuronal activations for individual concepts exhibit distinct, Gaussian-like distributions. This discovery allows for a more precise understanding of how neurons encode multiple concepts, enabling us to move beyond coarse, monolithic neuron-to-concept mappings. Building upon these insights, we proposed $\mathsf{NeuronLens}$, a novel range-based framework for neuron interpretation and manipulation. $\mathsf{NeuronLens}$ offers fine-grained control that reduces interference with unrelated concepts by attributing specific activation ranges within neurons to individual concepts. Extensive empirical evaluations demonstrated that $\mathsf{NeuronLens}$ outperforms existing methods in maintaining concept-specific precision while minimizing unintended side effects. The hyperparameter $\tau$ was set to ensure fair comparison across methods. Future work could explore mathematically deriving optimal $\tau$ values for individual neurons. Additionally, our work proved that the fundamental unit of interpretation can be a range within the activation space, which can be exploited for better interpretability.
8 Impact Statement
This work advances neural network interpretability by providing a fine-grained understanding of concept encoding in language models. The proposed $\mathsf{NeuronLens}$ framework enables precise control of model behavior, benefiting research in model safety and reliability. While this improved understanding could potentially be misused, the work’s theoretical nature and focus on interpretability methods makes immediate harmful applications unlikely.
Appendix A Limitations
While $\mathsf{NeuronLens}$ can disentangle polysemanticity to a degree using Gaussian Like Distribution, it is unable to completely disentangle concepts encoded in the polysemantic neurons, because there still is a significant overlap in the distributions of concepts in activations. Additionally, in this work, we use $\tau$ to be a fixed value of 2.5 to make the comparison of approaches fair, but $\tau$ selection can be optimized to be more sophisticated. We also get results primarily from the penultimate layer, and not the intermediate or earlier layers, however, we do give ablation and rationale for this choice in Appendix E
Appendix B Training Details
For BERT, DistilBERT, and Llama, we utilize pretrained models. Since BERT, and DistilBert are not inherently trained as a conversational agent, we use top-performing fine-tuned models from the Hugging Face repository. For the Llama model, few-shot prompt completion is employed to predict class labels. This involves providing a small number of training samples from the dataset to guide the model’s predictions.
For GPT-2, we fine-tune the pretrained model across all datasets for three epochs. The input sequence is constructed by concatenating the text with a <sep> token, followed by the class label, and ending with an <eos> token. During training, the loss is back-propagated only for the class label token, while all other tokens are assigned a skip label (-100). Additionally, all class labels are added to the model’s dictionary as special single-token entries.
Dataset Preprocessing for Llama For Llama we process whole datasets in few shout settings and only curate 2000 samples per class, where the model prediction was correct.
Appendix C Saliency details
Max activations. Frankle & Carbin (2019) extract high neural activations as a saliency ranking metric relying upon the rationale that maximally activating neurons are salient as these neurons play a critical role in controlling the model’s output, highlighting their importance for a concept $c$.To identify them, the column-wise mean of absolute neuronal activations in $H^{l}_{c}$, $H^{l}_{c}$ is defined in Section2.3, is computed, given that high negative activations also carry significant signals (Voita et al., 2023). The magnitude of the means is then considered as a ranking for concept $c$.
Probe analysis. Dalvi et al. (2019b) train a linear classifier on the hidden representations $H^{l}_{c}$ to distinguish between concepts. The learned model weights are then utilized as a saliency ranking. This process involves learning a weight matrix $W\in\mathbb{R}^{d\times|c|}$, where $d$ is the hidden dimension and $|c|$ is the number of concept classes. The absolute weight values of each row in the weight matrix are used as a ranking for the importance of each neuron for a given concept. To prevent the emergence of redundant solutions characterized by minimal variations in the weights, the probe is trained using the elastic regularization technique.
Probeless. Antverg & Belinkov (2022) examine individual neurons, without the need for auxiliary classifiers, using the element-wise difference between mean vectors. The element-wise difference between mean vectors is computed as $r=\sum_{c,c^{\prime}\in C}|q(c)-q(c^{\prime})|$, where $r\in\mathbb{R}^{d}$ and $d$ is the hidden dimension. The final neuron saliency ranking is obtained by sorting $r$ in descending order.
Appendix D Full Results
Here we provide the complete results for the datasets shown in Table 3. In Table 5 we provide results on IMDB dataset on all selected models. In Table 6 we provide results on SST2 dataset on all selected models. In Table 7 we provide results on Emotions dataset on all selected models. In Table 8 we provide results on DBPedia-14 dataset on all selected models.
Table 5: Evaluation of selected models on the IMDB dataset using neuron and range masking techniques. Here, Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance.
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
BERT Class 0 0.930 0.908 0.926 0.901 -0.169 -0.352 0.061 -0.066 -0.163 -0.359 0.059 0.035
Class 1 0.926 0.901 0.930 0.908 -0.211 -0.355 0.057 -0.091 -0.206 -0.361 0.056 0.025
GPT-2 Class 0 0.965 0.941 0.940 0.922 -0.935 -0.922 0.050 0.057 -0.905 -0.901 0.055 0.046
Class 1 0.940 0.922 0.965 0.941 -0.620 -0.667 0.005 0.018 -0.610 -0.657 0.015 0.027
Llama-3.2-3B Class 0 1.000 0.619 1.000 0.500 -0.643 -0.448 -0.515 -0.287 -0.640 -0.446 -0.502 -0.278
Class 1 1.000 0.500 1.000 0.619 -0.877 -0.410 -0.273 -0.304 -0.873 -0.409 -0.265 -0.303
Table 6: Evaluation of selected models on the SST2 dataset using neuron and range masking techniques. Here, Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance.
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
BERT Class 0 0.890 0.882 0.930 0.925 -0.058 -0.308 0.029 -0.047 -0.075 -0.329 0.031 0.036
Class 1 0.930 0.925 0.890 0.882 -0.043 -0.318 0.033 -0.045 -0.045 -0.330 0.030 0.050
GPT-2 Class 0 0.950 0.937 0.981 0.978 -0.142 -0.158 0.010 0.012 -0.142 -0.167 0.009 0.010
Class 1 0.981 0.978 0.950 0.937 -0.187 -0.223 0.041 0.053 -0.176 -0.216 0.041 0.046
Llama-3.2-3B Class 0 1.000 0.620 1.000 0.690 -0.532 -0.459 -0.420 -0.424 -0.532 -0.456 -0.404 -0.415
Class 1 1.000 0.690 1.000 0.620 -0.289 -0.379 -0.326 -0.315 -0.284 -0.376 -0.306 -0.301
Table 7: Evaluation of selected models on the Emotions dataset using neuron and range masking techniques. Here, Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance.
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
BERT Class 0 0.960 0.935 0.901 0.851 -0.241 -0.718 0.013 -0.266 -0.222 -0.718 0.012 -0.055
Class 1 0.942 0.904 0.905 0.861 -0.223 -0.691 0.028 -0.254 -0.213 -0.692 0.032 -0.064
Class 2 0.824 0.723 0.926 0.889 -0.371 -0.533 0.016 -0.284 -0.352 -0.534 0.018 -0.115
Class 3 0.927 0.873 0.916 0.876 -0.247 -0.664 0.010 -0.256 -0.240 -0.667 0.012 -0.057
Class 4 0.884 0.837 0.922 0.880 -0.406 -0.646 0.012 -0.251 -0.402 -0.648 0.012 -0.066
Class 5 0.591 0.566 0.929 0.886 -0.303 -0.392 0.004 -0.299 -0.303 -0.397 0.005 -0.090
GPT-2 Class 0 0.969 0.956 0.913 0.903 -0.695 -0.751 -0.125 -0.124 -0.698 -0.749 -0.009 -0.009
Class 1 0.939 0.938 0.925 0.908 -0.879 -0.882 -0.019 -0.009 -0.879 -0.880 -0.016 -0.015
Class 2 0.902 0.872 0.932 0.923 -0.776 -0.736 -0.029 -0.032 -0.780 -0.739 -0.023 -0.028
Class 3 0.910 0.905 0.932 0.921 -0.713 -0.714 -0.006 -0.007 -0.715 -0.716 -0.002 -0.001
Class 4 0.869 0.854 0.938 0.927 -0.754 -0.753 -0.240 -0.248 -0.754 -0.753 -0.127 -0.133
Class 5 0.857 0.798 0.932 0.923 -0.587 -0.601 -0.301 -0.308 -0.587 -0.601 -0.280 -0.289
Llama-3.2-3B Class 0 0.950 0.550 0.782 0.455 -0.950 -0.547 -0.655 -0.408 -0.945 -0.547 -0.571 -0.378
Class 1 0.905 0.498 0.804 0.473 -0.855 -0.495 -0.743 -0.433 -0.867 -0.494 -0.607 -0.404
Class 2 0.785 0.421 0.827 0.483 -0.785 -0.420 -0.771 -0.454 -0.785 -0.420 -0.658 -0.436
Class 3 0.790 0.482 0.833 0.476 -0.760 -0.477 -0.635 -0.423 -0.755 -0.476 -0.544 -0.402
Class 4 0.780 0.487 0.829 0.476 -0.780 -0.486 -0.534 -0.365 -0.780 -0.486 -0.444 -0.324
Class 5 0.536 0.296 0.855 0.498 -0.417 -0.284 -0.751 -0.465 -0.429 -0.282 -0.653 -0.434
Table 8: Evaluation of selected models on the DBPedia-14 dataset using neuron and range masking techniques. Here, Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance.
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
BERT Class 0 0.972 0.966 0.992 0.991 -0.082 -0.702 0.001 -0.014 -0.076 -0.698 0.001 -0.000
Class 1 0.987 0.986 0.991 0.990 -0.030 -0.778 $0$0.017 -0.018 -0.770 $0$0.000
Class 2 0.987 0.985 0.991 0.990 -0.239 -0.814 0.001 -0.018 -0.217 -0.806 0.001 -0.000
Class 3 0.997 0.997 0.990 0.989 -0.008 -0.766 $0$0.019 -0.001 -0.731 $0$0.000
Class 4 0.984 0.983 0.991 0.990 -0.058 -0.777 0.001 -0.018 -0.032 -0.761 $0$0.000
Class 5 0.995 0.995 0.990 0.989 -0.007 -0.795 $0$0.017 -0.001 -0.771 $0$0.000
Class 6 0.975 0.974 0.992 0.991 -0.121 -0.807 $0$0.015 -0.112 -0.803 $0$0.001
Class 7 0.994 0.994 0.990 0.989 -0.028 -0.789 $0$0.017 -0.010 -0.767 $0$0.000
Class 8 1.000 1.000 0.990 0.989 -0.001 -0.808 $0$0.022 $0$0.772 $0$0.000
Class 9 0.999 0.998 0.990 0.989 -0.004 -0.837 $0$0.019 -0.001 -0.811 $0$0.000
Class 10 0.994 0.993 0.990 0.989 -0.025 -0.846 $0$0.016 -0.005 -0.831 $0$0.000
Class 11 0.997 0.997 0.990 0.989 -0.013 -0.751 $0$0.017 -0.001 -0.726 $0$0.000
Class 12 0.990 0.990 0.990 0.989 -0.018 -0.772 $0$0.017 -0.005 -0.755 $0$0.000
Class 13 0.994 0.994 0.990 0.989 -0.009 -0.740 0.001 -0.017 -0.001 -0.721 $0$0.000
GPT-2 Class 0 0.985 0.977 0.990 0.989 -0.860 -0.877 -0.133 -0.136 -0.850 -0.869 -0.002 -0.017
Class 1 0.995 0.992 0.990 0.988 -0.500 -0.567 -0.180 -0.192 -0.460 -0.544 -0.023 -0.024
Class 2 0.985 0.980 0.990 0.989 -0.890 -0.904 -0.189 -0.213 -0.880 -0.902 -0.004 -0.010
Class 3 0.995 0.995 0.990 0.987 -0.900 -0.933 -0.145 -0.143 -0.900 -0.927 -0.008 -0.017
Class 4 0.970 0.969 0.992 0.989 -0.715 -0.773 -0.224 -0.260 -0.695 -0.750 -0.042 -0.062
Class 5 0.995 0.993 0.990 0.988 -0.315 -0.446 -0.127 -0.192 -0.290 -0.432 -0.013 -0.025
Class 6 0.965 0.964 0.992 0.990 -0.925 -0.932 -0.052 -0.062 -0.910 -0.928 -0.006 -0.007
Class 7 1.000 0.998 0.989 0.987 -0.815 -0.865 -0.003 -0.008 -0.775 -0.846 -0.026 -0.057
Class 8 1.000 1.000 0.989 0.987 -0.995 -0.990 -0.148 -0.188 -0.900 -0.932 -0.026 -0.055
Class 9 1.000 1.000 0.989 0.987 -0.975 -0.979 -0.250 -0.268 -0.955 -0.958 -0.020 -0.049
Class 10 0.995 0.993 0.990 0.988 -0.595 -0.685 -0.045 -0.053 -0.590 -0.675 -0.005 -0.011
Class 11 0.985 0.984 0.990 0.988 -0.210 -0.453 -0.094 -0.118 -0.135 -0.396 -0.015 -0.034
Class 12 0.990 0.988 0.990 0.988 -0.930 -0.938 -0.293 -0.309 -0.855 -0.880 -0.013 -0.029
Class 13 1.000 0.999 0.989 0.987 -0.985 -0.986 -0.393 -0.416 -0.945 -0.981 -0.018 -0.044
Llama-3.2-3B Class 0 1.000 0.586 1.000 0.559 -0.990 -0.584 -0.949 -0.473 -0.990 -0.584 -0.823 -0.441
Class 1 1.000 0.533 1.000 0.563 -1.000 -0.528 -0.870 -0.446 -0.970 -0.528 -0.706 -0.371
Class 2 1.000 0.467 1.000 0.568 -0.995 -0.462 -0.963 -0.477 -0.995 -0.461 -0.838 -0.432
Class 3 1.000 0.460 1.000 0.569 -0.995 -0.459 -0.981 -0.486 -0.995 -0.459 -0.815 -0.420
Class 4 1.000 0.828 1.000 0.539 -0.965 -0.809 -0.981 -0.454 -0.955 -0.808 -0.852 -0.412
Class 5 1.000 0.349 1.000 0.568 -1.000 -0.348 -0.882 -0.429 -0.989 -0.347 -0.585 -0.346
Class 6 1.000 0.809 1.000 0.541 -1.000 -0.787 -0.972 -0.449 -1.000 -0.787 -0.736 -0.366
Class 7 1.000 0.599 1.000 0.558 -0.855 -0.588 -0.918 -0.410 -0.860 -0.586 -0.489 -0.274
Class 8 1.000 0.420 1.000 0.572 -1.000 -0.420 -0.957 -0.467 -1.000 -0.420 -0.660 -0.335
Class 9 1.000 0.527 1.000 0.563 -1.000 -0.524 -0.842 -0.435 -0.995 -0.523 -0.552 -0.320
Class 10 1.000 0.505 1.000 0.565 -0.995 -0.503 -0.907 -0.464 -1.000 -0.503 -0.589 -0.322
Class 11 1.000 0.505 1.000 0.565 -0.975 -0.501 -0.862 -0.416 -0.970 -0.501 -0.579 -0.313
Class 12 1.000 0.560 1.000 0.561 -0.980 -0.545 -0.812 -0.417 -0.975 -0.544 -0.496 -0.310
Class 13 1.000 0.587 1.000 0.559 -0.990 -0.584 -0.722 -0.406 -0.985 -0.584 -0.588 -0.337
Appendix E Layer Ablation
In Table 9 and Table 10 we provide results of applying both approaches on all layers of GPT-2 model on Emotions dataset. From the results we can see that: Early layers (1-3) show highly variable and often severe impacts: Layer 1 exhibits minimal effects ($\Delta Acc=-0.113$, $\Delta CAcc=-0.064$), while Layers 2-3 show extreme degradation ($\Delta Acc\approx-0.7$, $\Delta CAcc>-0.5$). Middle layers (4-8) demonstrate inconsistent behavior with high variance in impacts. Layer 12, however, achieves an optimal balance: it maintains substantial primary task impact ($\Delta Acc=-0.571$) while minimizing auxiliary concept interference ($\Delta CAcc=-0.060$). This pattern holds true for both neuron masking and range masking techniques, with range masking showing slightly better preservation of auxiliary concepts ($\Delta CAcc=-0.045$). The mid-range primary task degradation combined with minimal auxiliary impact makes Layer 12 the most suitable for targeted interventions, offering better control and specificity compared to earlier layers.
Table 9: Evaluation of layer selection on GPT-2 model on the Emotions dataset using neuron and range masking techniques. 20% Neurons selected. Here, Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance.
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
1 Class 0 0.970 0.957 0.915 0.904 -0.029 -0.071 -0.074 -0.100 0.006 0.002 -0.004 -0.005
Class 1 0.933 0.932 0.931 0.913 -0.011 -0.056 -0.090 -0.116 0.001 -0.003 -0.004 -0.004
Class 2 0.901 0.865 0.934 0.924 -0.206 -0.195 -0.052 -0.092 -0.019 -0.015 -0.001 -0.002
Class 3 0.926 0.924 0.932 0.919 -0.128 -0.152 -0.051 -0.090 -0.004 -0.005 -0.001 -0.002
Class 4 0.885 0.867 0.938 0.927 -0.055 -0.084 -0.061 -0.093 -0.016 -0.009 0.002 -0.001
Class 5 0.851 0.786 0.934 0.924 -0.249 -0.217 -0.055 -0.094 0.016 0.013 -0.004 -0.005
2 Class 0 0.970 0.957 0.915 0.904 -0.804 -0.808 -0.389 -0.386 -0.061 -0.133 -0.077 -0.096
Class 1 0.933 0.932 0.931 0.913 0.053 -0.003 -0.819 -0.781 -0.011 -0.049 -0.110 -0.145
Class 2 0.901 0.865 0.934 0.924 -0.868 -0.737 -0.515 -0.519 -0.365 -0.337 -0.077 -0.126
Class 3 0.926 0.924 0.932 0.919 -0.870 -0.805 -0.498 -0.501 -0.215 -0.248 -0.096 -0.153
Class 4 0.885 0.867 0.938 0.927 -0.729 -0.707 -0.461 -0.463 -0.042 -0.077 -0.076 -0.116
Class 5 0.851 0.786 0.934 0.924 -0.845 -0.769 -0.511 -0.508 -0.229 -0.188 -0.106 -0.163
3 Class 0 0.970 0.957 0.915 0.904 -0.896 -0.904 -0.824 -0.832 -0.647 -0.688 -0.517 -0.544
Class 1 0.933 0.932 0.931 0.913 -0.901 -0.916 -0.835 -0.832 -0.568 -0.607 -0.609 -0.630
Class 2 0.901 0.865 0.934 0.924 -0.868 -0.845 -0.838 -0.851 -0.605 -0.600 -0.589 -0.619
Class 3 0.926 0.924 0.932 0.919 -0.868 -0.896 -0.830 -0.840 -0.567 -0.605 -0.567 -0.596
Class 4 0.885 0.867 0.938 0.927 -0.800 -0.811 -0.849 -0.857 -0.502 -0.522 -0.513 -0.544
Class 5 0.851 0.786 0.934 0.924 0.022 0.081 -0.865 -0.881 -0.155 -0.124 -0.561 -0.596
4 Class 0 0.970 0.957 0.915 0.904 -0.650 -0.703 -0.698 -0.764 -0.608 -0.621 -0.499 -0.510
Class 1 0.933 0.932 0.931 0.913 -0.845 -0.884 -0.667 -0.725 -0.491 -0.519 -0.480 -0.491
Class 2 0.901 0.865 0.934 0.924 -0.858 -0.824 -0.772 -0.809 -0.488 -0.497 -0.506 -0.523
Class 3 0.926 0.924 0.932 0.919 -0.700 -0.808 -0.663 -0.739 -0.534 -0.546 -0.512 -0.528
Class 4 0.885 0.867 0.938 0.927 -0.239 -0.514 -0.754 -0.797 -0.304 -0.307 -0.452 -0.471
Class 5 0.851 0.786 0.934 0.924 -0.612 -0.463 -0.692 -0.765 -0.047 -0.038 -0.525 -0.541
5 Class 0 0.970 0.957 0.915 0.904 -0.838 -0.852 -0.492 -0.630 -0.695 -0.688 -0.554 -0.555
Class 1 0.933 0.932 0.931 0.913 -0.387 -0.563 -0.683 -0.714 -0.552 -0.564 -0.605 -0.599
Class 2 0.901 0.865 0.934 0.924 -0.702 -0.700 -0.634 -0.690 -0.472 -0.470 -0.607 -0.605
Class 3 0.926 0.924 0.932 0.919 -0.361 -0.507 -0.615 -0.692 -0.567 -0.575 -0.538 -0.539
Class 4 0.885 0.867 0.938 0.927 -0.873 -0.844 -0.525 -0.650 -0.668 -0.653 -0.594 -0.594
Class 5 0.851 0.786 0.934 0.924 -0.637 -0.573 -0.588 -0.681 -0.069 -0.022 -0.548 -0.553
6 Class 0 0.970 0.957 0.915 0.904 -0.720 -0.775 -0.829 -0.830 -0.484 -0.499 -0.318 -0.322
Class 1 0.933 0.932 0.931 0.913 -0.871 -0.887 -0.750 -0.768 -0.176 -0.195 -0.499 -0.499
Class 2 0.901 0.865 0.934 0.924 -0.895 -0.860 -0.735 -0.773 -0.680 -0.638 -0.335 -0.348
Class 3 0.926 0.924 0.932 0.919 -0.863 -0.884 -0.772 -0.793 -0.418 -0.431 -0.379 -0.381
Class 4 0.885 0.867 0.938 0.927 -0.621 -0.669 -0.743 -0.784 -0.430 -0.435 -0.247 -0.262
Class 5 0.851 0.786 0.934 0.924 -0.143 -0.086 -0.808 -0.831 -0.114 -0.070 -0.474 -0.478
Table 10: Evaluation of layer selection on GPT-2 model on the Emotions dataset using neuron and range masking techniques. 20% Neurons selected. Here, Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance.
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
7 Class 0 0.970 0.957 0.915 0.904 -0.908 -0.901 -0.752 -0.753 -0.527 -0.538 -0.492 -0.498
Class 1 0.933 0.932 0.931 0.913 -0.884 -0.895 -0.743 -0.729 -0.484 -0.509 -0.330 -0.338
Class 2 0.901 0.865 0.934 0.924 -0.866 -0.835 -0.767 -0.765 -0.451 -0.442 -0.336 -0.355
Class 3 0.926 0.924 0.932 0.919 -0.786 -0.819 -0.641 -0.666 -0.445 -0.457 -0.331 -0.346
Class 4 0.885 0.867 0.938 0.927 -0.626 -0.618 -0.810 -0.817 -0.341 -0.335 -0.521 -0.532
Class 5 0.851 0.786 0.934 0.924 0.106 0.147 -0.810 -0.811 0.102 0.107 -0.547 -0.553
8 Class 0 0.970 0.957 0.915 0.904 -0.776 -0.791 -0.209 -0.291 -0.191 -0.312 -0.082 -0.114
Class 1 0.933 0.932 0.931 0.913 -0.585 -0.667 -0.412 -0.441 -0.591 -0.644 -0.199 -0.227
Class 2 0.901 0.865 0.934 0.924 -0.692 -0.716 -0.469 -0.496 -0.560 -0.562 -0.468 -0.486
Class 3 0.926 0.924 0.932 0.919 -0.657 -0.714 -0.415 -0.464 -0.468 -0.503 -0.230 -0.266
Class 4 0.885 0.867 0.938 0.927 -0.501 -0.509 -0.531 -0.569 -0.201 -0.234 -0.258 -0.290
Class 5 0.851 0.786 0.934 0.924 -0.092 -0.050 -0.634 -0.647 0.065 0.058 -0.279 -0.308
9 Class 0 0.970 0.957 0.915 0.904 -0.759 -0.768 -0.311 -0.351 -0.610 -0.661 -0.307 -0.328
Class 1 0.933 0.932 0.931 0.913 -0.570 -0.713 -0.319 -0.346 -0.906 -0.910 -0.267 -0.298
Class 2 0.901 0.865 0.934 0.924 -0.424 -0.520 -0.504 -0.531 -0.635 -0.643 -0.579 -0.595
Class 3 0.926 0.924 0.932 0.919 -0.810 -0.834 -0.501 -0.502 -0.759 -0.772 -0.502 -0.516
Class 4 0.885 0.867 0.938 0.927 -0.358 -0.357 -0.476 -0.481 -0.587 -0.566 -0.519 -0.527
Class 5 0.851 0.786 0.934 0.924 -0.133 -0.101 -0.546 -0.554 0.106 0.104 -0.450 -0.462
10 Class 0 0.970 0.957 0.915 0.904 -0.733 -0.741 -0.105 -0.126 -0.624 -0.659 -0.146 -0.163
Class 1 0.933 0.932 0.931 0.913 -0.389 -0.671 -0.178 -0.209 -0.899 -0.911 -0.254 -0.285
Class 2 0.901 0.865 0.934 0.924 -0.230 -0.513 -0.116 -0.224 -0.699 -0.735 -0.409 -0.451
Class 3 0.926 0.924 0.932 0.919 -0.434 -0.687 -0.081 -0.133 -0.898 -0.905 -0.401 -0.455
Class 4 0.885 0.867 0.938 0.927 -0.489 -0.506 -0.188 -0.256 -0.140 -0.186 -0.063 -0.102
Class 5 0.851 0.786 0.934 0.924 -0.306 -0.243 -0.157 -0.240 0.063 0.010 -0.095 -0.127
11 Class 0 0.970 0.957 0.915 0.904 -0.358 -0.496 -0.382 -0.414 -0.301 -0.441 -0.121 -0.148
Class 1 0.933 0.932 0.931 0.913 -0.800 -0.857 -0.078 -0.123 -0.858 -0.875 -0.128 -0.162
Class 2 0.901 0.865 0.934 0.924 -0.897 -0.861 -0.416 -0.450 -0.901 -0.864 -0.464 -0.500
Class 3 0.926 0.924 0.932 0.919 -0.923 -0.921 -0.427 -0.470 -0.913 -0.914 -0.354 -0.393
Class 4 0.885 0.867 0.938 0.927 -0.152 -0.212 -0.039 -0.075 -0.210 -0.239 -0.181 -0.204
Class 5 0.851 0.786 0.934 0.924 0.047 -0.028 -0.131 -0.173 0.053 0.002 -0.142 -0.159
12 Class 0 0.970 0.957 0.915 0.904 -0.550 -0.603 -0.013 -0.003 -0.542 -0.594 0.005 0.012
Class 1 0.933 0.932 0.931 0.913 -0.526 -0.545 0.001 0.012 -0.521 -0.538 -0.005 -0.004
Class 2 0.901 0.865 0.934 0.924 -0.416 -0.402 0.002 0.006 -0.419 -0.407 0.007 0.006
Class 3 0.926 0.924 0.932 0.919 -0.561 -0.576 -0.007 0.003 -0.561 -0.572 0.000 0.005
Class 4 0.885 0.867 0.938 0.927 -0.655 -0.658 -0.042 -0.034 -0.657 -0.659 -0.011 -0.003
Class 5 0.851 0.786 0.934 0.924 -0.718 -0.672 -0.300 -0.297 -0.718 -0.672 -0.267 -0.266