### v2

# Neurons Speak in Ranges: Breaking Free from Discrete Neuronal Attribution

Abstract
Interpreting the internal mechanisms of large language models (LLMs) is crucial for improving their trustworthiness and utility. Prior work has primarily focused on mapping individual neurons to discrete semantic concepts. However, such mappings struggle to handle the inherent polysemanticity in LLMs, where individual neurons encode multiple, distinct concepts. Through a comprehensive analysis of both encoder and decoder-based LLMs across diverse datasets, we observe that even highly salient neurons, identified via various attribution techniques for specific semantic concepts, consistently exhibit polysemantic behavior. Importantly, activation magnitudes for fine-grained concepts follow distinct, often Gaussian-like distributions with minimal overlap. This observation motivates a shift from neuron attribution to range-based interpretation. We hypothesize that interpreting and manipulating neuron activation ranges would enable more precise interpretability and targeted interventions in LLMs. To validate our hypothesis, we introduce $\mathsf{NeuronLens}$, a novel range-based interpretation and manipulation framework that provides a finer view of neuron activation distributions to localize concept attribution within a neuron. Extensive empirical evaluations demonstrate that $\mathsf{NeuronLens}$ significantly reduces unintended interference, maintaining precise manipulation of targeted concepts, outperforming neuron attribution.
1 Introduction
Large language models (LLMs) demonstrate remarkable performance across natural language understanding, generation, and transformation tasks (Brown et al., 2020; Bommasani et al., 2022; Touvron et al., 2023; Raffel et al., 2019). However, the inner workings of LLMs remain largely opaque (Burkart and Huber, 2021), as their representations are distributed across billions of parameters. This lack of interpretability raises critical concerns about reliability, fairness, and trustworthiness, particularly in high-stakes domains such as healthcare, law, and education. To this end, neuron-level interpretability can address these concerns by enabling researchers to uncover how individual neurons encode semantic concepts and contribute to model outputs. With this understanding, researchers can diagnose safety risks (Wei et al., 2024; He et al., 2024), refine model outputs (Meng et al., 2023; Rizwan et al., 2024), optimize efficiency through pruning (Frankle and Carbin, 2019; Haider and Taj, 2021), and steer model’s representations toward desired objectives (Subramani et al., 2022; Li et al., 2023; Rodriguez et al., 2024).
Recent research efforts have made significant progress in neuron-level interpretation by identifying salient neurons that influence model behavior (Dalvi et al., 2019a; Antverg and Belinkov, 2022; Conmy et al., 2023; Marks et al., 2024). Approaches such as maximal activation analysis (Foote et al., 2023; Frankle and Carbin, 2019), Probe-based methods that employ auxiliary classifiers to distinguish between concepts (Dalvi et al., 2019a, b), and the probeless approach bypasses the need for classifiers by directly analyzing neurons (Antverg and Belinkov, 2022). Other techniques include circuit discovery, attributes concepts to groups of interacting neurons (Olah et al., 2020; Conmy et al., 2023), and causal analysis, identify the internal components role in model behavior (Vig et al., 2020).
Figure 1: $\mathsf{NeuronLens}$ leverages distinct, Gaussian-like activation patterns to enable fine-grained concept attribution.
While these methods have advanced our understanding of neurons, they often rely on discrete neuron-to-concept mappings, which assume that entire neurons encode single concepts. However, neurons frequently exhibit polysemanticity; the ability to encode multiple, seemingly unrelated concepts (Lecomte et al., 2024; Marshall and Kirchner, 2024). Given this heterogeneous encoding of concepts, traditional approaches often lead to unintended consequences when manipulating neurons, as changes intended for one concept may inadvertently affect others encoded by the same neuron or suboptimal interpretations of concepts (Sajjad et al., 2022).
Despite being traditionally viewed as a challenge, could polysemanticity instead provide a unique lens for advancing interpretability and model control? If individual neurons encode multiple concepts, might their activation spectrum reveal distinct and identifiable patterns for each concept? Could these patterns enable precise interventions that adjust one concept while minimizing interference with others, overcoming the limitations of coarse, monolithic neuron-to-concept mappings?
This work seeks to address these questions by analyzing the activation patterns of neurons in both encoder-based and decoder-based LLMs. Through statistical and qualitative analysis, we find that neuronal activations for concepts follow Gaussian-like distributions, with distinct patterns for different concepts. Our key insight is that the unit of interpretability lies at a level more fine-grained than the neuron itself. Within a neuron’s activation spectrum, activation ranges corresponding to specific concepts can be used as a finer unit of interpretability. This nuanced perspective enables a more precise approach to neuron interpretation and manipulation, addressing the limitations of traditional, discrete neuron-to-concept mappings.
Building upon these insights, we introduce $\mathsf{NeuronLens}$ visualised in Figure 1, a range-based framework for neuron interpretation and manipulation. $\mathsf{NeuronLens}$ identifies and maps specific activation ranges within a neuron’s distribution to individual concepts, rather than attributing entire neurons to single concepts. For each concept, $\mathsf{NeuronLens}$ calculates a range that covers its activation spectrum, capturing the concept-specific activations while excluding unrelated concepts. Through experiments on encoder-based and decoder-based LLMs across several text classification datasets, we show that $\mathsf{NeuronLens}$ significantly reduces unintended interference by up to 25 percentage points in auxiliary concepts and up to 7x in LLM latent capabilities, while maintaining precise manipulation of targeted concepts, outperforming existing methods.
Our key contributions are: (1) To the best of our knowledge, this is the first work that performs a comprehensive study unfolding polysemantic neurons using activation spectrums. (2) We show that neuronal activations in LLMs form distinct concept level Gaussian-like distributions, with salient neurons exhibiting limited overlap in their activation patterns across concepts. (3) We empirically demonstrate that activation ranges within a neuron’s activation spectrum offer a more precise unit of interpretability, offering a refined framework for neuron-level analysis. (4) We propose $\mathsf{NeuronLens}$, an activation range-based framework for interpreting and manipulating neuronal activations, which enables fine-grained concept attribution and reduces unintended interference compared to neuron level intervention.
2 Neuron Interpretation Analysis
This section provides an overview of the neuron analysis, methods for extracting salient neurons, and causally validating their saliency.
2.1 Preliminaries
Neuron. We refer to the output of an activation as a neuron. In a transformer model, we consider neurons of hidden state vectors of different transformer layers. Formally, given a hidden state vector $\mathbf{h}^{l}\in\mathbb{R}^{d}$ of size $d$ produced by layer $l$, $h^{l}_{j}$ denotes its $j$-th neuron, i.e., the $j$-th component of $\mathbf{h}^{l}$.
Concept. A concept $c\in C$ is a high-level semantic category that groups each input instance (or components of every instance), where $C$ is the set of all concepts. For example, in a language task, a sentence can be categorized into 4 types: declarative, interrogative, imperative, and exclamatory, where each type is a concept. Words of a sentence can also have concepts like nouns, verbs, adjectives, adverbs, etc. In this study, we focus on the situation where all input samples are labelled with concepts.
Saliency Ranking. A saliency ranking orders the importance of neurons based on some saliency metric. For a hidden state vector $\mathbf{h}^{l}\in\mathbb{R}^{d}$, $s_{j,c}$ denotes the value of the saliency metric for the $j$-th neuron with respect to a concept $c$. The saliency ranking $(r_{c}(1),r_{c}(2),\cdots,r_{c}(d))$ is a permutation of the indices of neurons $(1,2,\cdots,d)$, where $r_{c}(j)<r_{c}(i)$ if $s_{j,c}>s_{i,c}$. The saliency metric is usually predetermined, e.g., the absolute value of each neuron.
Concept Learning. Given a hidden state vector $\mathbf{h}^{l}$ as input, the associated concept can be the output of an appended neural network (e.g., several fully connected layers). The parameters of this appended neural network can be trained using training samples labelled with concepts.
Metrics. To establish the causal validity of the attribution, we employ two quantitative metrics: prediction accuracy and the model’s predictive probability as a proxy for confidence score. First, baseline measurements of both accuracy and confidence for all concepts $C$ without any intervention (unmodified model) are established. Post-intervention measurements are recorded for the target concept $c$ and auxiliary concepts(other concepts in the dataset) $c^{\prime}$. The effectiveness and precision of attribution are assessed through two key metrics: (1) the magnitude of performance degradation for concept $c$, and (2) the extent of unintended impact on auxiliary concepts $c^{\prime}$. Throughout our analysis, we denote the accuracy and confidence metrics for concept $c$ as Acc and Conf respectively, while corresponding measurements for auxiliary concepts $c^{\prime}$ are represented as CAcc and CConf. For evaluating the effect of the interventions on LLMs latent capabilities, we utilize perplexity (PPL) and MMLU (Hendrycks et al., 2021) zero-shot accuracy.
Table 1: Performance drops relative to Baseline configuration (i.e., unaltered model’s performance) for three techniques: Probeless, Probe, and Max. All values show the difference from Base values. Results are for Emotions dataset on the GPT-2 model using 30% salient neurons of each method. Metrics are detailed in Section 2.1.
Probeless Probe Max
 Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
-0.524 -0.510 -0.086 -0.086 -0.052 -0.036 -0.018 -0.049 -0.735 -0.739 -0.103 -0.103
2.2 Concept Erasure
To assess the performance of a neuronal attribution, concept erasure acts as a critical diagnostic intervention to determine the causal effect of identified neurons for a given concept (Dalvi et al., 2019b; Dai et al., 2022; Dalvi et al., 2019c; Morcos et al., 2018). The core idea is that if a neuron is salient to a concept, eliminating it should result in the degradation of that concept’s performance while causing minimal disruption to other concepts. This can be formalized as follows: given a concept-learning model $M$ that maps any input instance $x$ (or part of an instance) to a concept $M(x)=c\quad c\in C$, an ideal intervened model $M^{\prime}_{\text{ideal}}$ after erasing a target concept $c\in C$ should satisfy the following property:
$M^{\prime}_{\text{ideal}}(x)=\begin{cases}\neq M(x)&\text{if }M(x)=c,\\ =M(x)&\text{if }M(x)\neq c.\end{cases}$
A popular approach of concept erasure in neuronal analysis literature (Dai et al., 2022; Antverg and Belinkov, 2022) is zeroing out specific neurons that are “important” to the target concept. Other studies have argued that zeroing out neurons is an overly aggressive intervention that can lead to catastrophic degradation in model performance. In Appendix Section D, we provide an ablation comparing different activation interventions for concept erasure.
2.3 Salient Neurons Extraction
Problem setup and preparation: We record activations for training samples of different concepts to perform neuron interpretation. Specifically, if we want to interpret neurons of $\mathbf{h}^{l}$ (hidden state vector at layer $l$), we traverse the training dataset and store the values of $\mathbf{h}^{l}$ and the associated concepts of all samples into a set $H^{l}$. The set $H^{l}$ is further partitioned into $H^{l}_{c}$ for all concepts $c\in C$. Such preparation is common in the relevant literature (Dalvi et al., 2019c, b; Antverg and Belinkov, 2022).
Max activations. Frankle and Carbin (2019) extract high activations as a saliency ranking metric relying upon the rationale that maximally activating neurons with respect to a concept $c$ are important for that concept. Probe analysis. Dalvi et al. (2019b) train a linear classifier on the hidden representations $H^{l}_{c}$ to predict each concept. The learned model weights are then utilized as a saliency ranking. Probeless. Antverg and Belinkov (2022) examine individual neurons, without the need for auxiliary classifiers, using the element-wise difference between mean vectors. Details of these approaches are provided in Appendix G.
To assess the effectiveness of the attribution methods, we perform neuron masking on the salient neurons identified by each method in a concept erasure task. Table 1 provides the results for this experiment. We observe that irrespective of the method used to obtain saliency ranking, a single concept eraser using salient neurons causes deterioration in performance across several concepts. Max activation causes the highest degree of deterioration in the targeted concept while maintaining a comparable deterioration in auxiliary concepts. Based on this finding, we adopt max activation ranking for saliency ranking. Moreover, we hypothesize that one reason for such deterioration in overall performance is due to the polysemantic nature of neurons.
Figure 2: Overlap of top 30% salient neurons across classes.
3 Polysemanticity
The polysemanticity of neuronal units, including salient neurons that encode information about multiple concepts, poses a challenge to neural network interpretation and manipulation. In this section, we discuss the degree of polysemanticity in salient neurons in detail.
Polysemanticity often arises when models must represent more features than their capacity allows or due to specific training paradigms. Limited representational space forces neurons to encode multiple unrelated features to maintain performance (Anthropic, 2023). Training methods like subword tokenization, designed to reduce vocabulary size and model complexity, lead to context-dependent token splits, causing neural activations to encode multiple meanings (Sennrich et al., 2016; Elhage et al., 2022; Meng et al., 2022). Additionally, Lecomte et al. (2024) show that even with sufficient capacity, certain weight initializations can induce polysemanticity by placing neurons near multiple conceptual regions.
Polysemanticity in salient neurons. Given that salient neurons have a strong causal association with the concept of interest, their tendency to be mono-semantic should be is high, but we find that there is a high degree of polysemanticity in salient neurons. We investigate this by extracting 30% salient neurons (i.e.: Max neurons) for different datasets on the GPT-2 model. The results show that there is a considerable overlap of salient neurons between concepts (classes) as shown in Figure 2. In the case of a two-class dataset IMDB, the overlap of salient neurons, selected by max, is more than 60%. This shows a high degree of polysemanticity. Consequently, we extrapolate that salient neural representations may exist in a polysemantic configuration, wherein a subset of the salient neurons encode information through intricate activation patterns.
The monolithic attribution paradigm potentially oversimplifies the complex, distributed nature of neuronal activation as can be seen in polysemanticity (Lecomte et al., 2024; Marshall and Kirchner, 2024) where a single neuron learns multiple seemingly unrelated concepts and elucidates them at different activation values.
4 Neuronal Activation Patterns
In this section, we analyze the properties of neuronal activations of the salient neurons (including polysemantic) extracted via maximal activation. Similar to (Gurnee et al., 2024), our findings indicate that neuronal activations form a Gaussian-like distribution. We further find that salient neurons have a distinct Gaussian distribution of activations for different concepts with limited overlap with other concept activations.
Figure 3: Neuronal Activation Patterns of six neurons on AG-News dataset class 1. Neurons 418 and 447 are the highest activating neurons, neurons 132 and 387 are middle-ranked neurons, and neurons 721 and 365 are the lowest activating neurons.
Figure 4: Comparison of neurons 480 and 675 showing class-specific activation patterns and fitted Gaussian curves. Both neurons were salient across all classes in top 5% on AG-News.
Qualitative Evaluation. To visually demonstrate that neuron activations for a concept $c$ follows a Gaussian-like distribution, we extract model representations as described in Section 2.1. Using saliency ranking $r_{c}$ for a single concept class, we examine neurons from different ranking positions in the GPT-2 model on the AG-News dataset: two top-ranked neurons ($r_{c}\leq 2$), two middle-ranked neurons ($r_{c}\approx d/2$), and two bottom-ranked neurons ($r_{c}\geq d-1$). In Figures 4 and 4, we use Kernel Density Estimation (KDE) to visualize these distributions. Figure 4 reveals that while the activations are Gaussian-like for different concepts, salient neurons demonstrate distinct activation patterns with limited overlap, middle-ranked neurons show a higher degree of overlap than the top ones, whereas non-salient neurons (bottom two) exhibit the highest overlap in their activation distributions.
Additionally in Figure 4, we identify and visualize two distinct types of polysemantic neurons that appear in the salient sets across all classes, when 5% salient set was selected, in the dataset. The first type, exemplified by neuron 480, maintains partially separable activation patterns despite being polysemantic, suggesting some degree of class-specific behavior. In contrast, the second type, represented by neuron 675, exhibits completely overlapping activation patterns across all classes, making it hard to disentangle. To further investigate this phenomenon, Figure 5 presents a broader analysis of neurons from the polysemantic subset, identified using a 5% saliency threshold(top 5% salient neurons selected for a concept). By examining these neurons’ behavior across four randomly selected classes (out of 14 total classes), we observe that most polysemantic neurons exhibit a high degree of separability, for some classes, while they respond to multiple classes, they tend to operate in partially separable activation ranges, supporting the possibility of meaningful disentanglement.
Figure 5: Box plot of neural activation of 11 polysemantic neurons (i.e: neurons in the salient group for all classes, percentage selected: 5% top salient) for 4 randomly selected classes out of 14 classes of DBPedia-14 dataset.
Quantitative Evaluation. To quantify the effect of Gaussian-like distribution of neurons for $c\in C$, we perform statistical analysis of activations. We computed the skewness, kurtosis (Joanes and Gill, 1998) and analyzed the normality of neuronal activations using Kolmogorov-Smirnov (KS) test (Massey Jr, 1951). Table 2 presents the results for distributional properties across all neurons. The average skewness is close to 0 across all datasets, indicating strong symmetry (ideal normal distribution: 0), and the average kurtosis is close to 3, nearly identical to the expected value for a normal distribution (3.0).
Table 2: Skewness, kurtosis, and Kolmogorov-Smirnov test results across various datasets.GPT-2 model
Dataset Skewness Kurtosis KS-Test
stanfordnlp/imdb 0.0014 3.6639 1.0000
fancyzhx/dbpedia_14 -0.0007 3.9360 0.9782
dair-ai/emotion 0.0015 3.0198 0.9446
fancyzhx/ag_news -0.0013 3.2060 0.9918
stanfordnlp/sst2 -0.0083 3.2038 1.0000
To quantitatively assess normality, while accounting for practical significance, we employ the KS test with an effect size threshold of 10%. This approach tests whether the distribution remains within a reasonable bound of normality, rather than testing for perfect normality, which is overly strict for real-world data. For each neuron, we normalize the activations to zero mean and unit variance, then compute the KS statistic against a standard normal distribution. The KS statistic represents the maximum absolute difference between the empirical and theoretical cumulative distribution functions. Using a threshold of 0.1 (allowing a maximum 10% deviation from normal), we find that close to 100% of the neurons exhibit practically normal distributions. The combination of near-ideal skewness and kurtosis values, visual confirmation through KDEs, and our effect size-based KS tests provide strong evidence that the activations follow approximately normal distributions.
We report quantitative statistics for all layers in Appendix Section I.1, which show that as layer depth increases, kurtosis steadily converges toward the Gaussian benchmark of 3.0, skewness remains near zero, and the 10% practical-normality score stays close to 1 across the network. A qualitative, layer-wise examination in Appendix Section I.2 further reveals that while all layers exhibit class level Gaussian-like activations. Early layers show substantial overlap between classes, this is consistent with the understanding that earlier layers focus on low-level features, not high level features like class. Beginning as early as layers 5–6, distinct class-specific Gaussians emerge and become progressively more separable in deeper layers, indicating a transition toward higher-level semantic representations.
5 Activation Ranges-guided Concept Erasure
Given that neuronal activations exhibit approximately Gaussian-like distributions with separable means, we can interpret and intervene on neurons more precisely than by ablating entire units. Specifically, $\mathsf{NeuronLens}$ ablates salient neurons identified through saliency ranking only when their activation falls within a selected range. The key idea is to identify a range that is strongly associated with the target concept $c$ intended for erasure. This range-based approach enables fine-grained ablation, thereby reducing unintended interference with non-target concepts. To validate our approach, we evaluate the causal efficacy of our method relative to neuron ablation using concept erasure experiments and assess the model’s latent capabilities following this intervention.
To calculate the aforementioned range, the framework utilises the means and standard deviations of the neuron activations. Specifically, first the empirical average $\mu\in\mathds{R}$ and standard deviation $\sigma\geq 0$ of the values of the salient neuron for all samples associated with the target concept $c\in C$ are calculated. After that, range is assigned as $[\mu-\tau\times\sigma,\ \mu+\tau\times\sigma]$, where $\tau>0$ is a hyperparameter to make a tradeoff between erasing the target concept $c$ (using larger $\tau$) and smaller impact on auxiliary concepts and general LLM capabilities (using smaller $\tau$). For this work, $\tau$ is set to $\tau=2.5$, assuming a fully Gaussian distribution. This threshold corresponds to a coverage of approximately $98.76\%$ of the distribution’s values, providing a slightly conservative bound for range-based interventions. Ablations for varying the hyperparameter $\tau$ are presented in Appendix Section H, the results indicate that targeted concept deteriorates up to 2.4-2.7 $\tau$ then plateaus, while auxiliary concepts begin to degrade further.
$h_{j}^{l}(x)=\begin{cases}\phi(x)&\text{if }h_{j}^{l}(x)\in\text{CR}(l,j,c)\\ h_{j}^{l}(x)&\text{otherwise}\end{cases}$
$\text{CR}(l,j,c)=[\mu\!-\!2.5\sigma,\ \mu\!+\!2.5\sigma],$
$\mu=\frac{1}{|H^{l}_{c}|}\sum_{\mathbf{h}^{l}\in H^{l}_{c}}h_{j}^{l},\sigma=\sqrt{\frac{1}{|H^{l}_{c}|}\sum_{\mathbf{h}^{l}\in H^{l}_{c}}(h_{j}^{l}-\mu)^{2}}$
where CR represents Correlated Range and $\phi()$ is the activation intervention function, which returns zero for the results presented in the main paper.
Notice that $H^{l}_{c}$ was defined in problem setup and preparation of section 2.3, which denotes the set of hidden state vector $\mathbf{h}^{l}(x_{c})$ at layer $l$ for all training samples $x_{c}$ associated with concept $c$. Here ||$\cdot$|| ⋅ | denotes the cardinality of a set.
5.1 Experimental Setup
Models. This study employs both encoder and decoder-based models, including fine-tuned BERT (Devlin et al., 2019), DistilBERT (Sanh et al., 2020), GPT-2 (Radford et al., 2019), and pretrained Llama-3.2-3B (Grattafiori, 2024). We incorporate our methodology at the penultimate layer; ablation for layer selection is provided in the Appendix Section I. The training details for the models are provided in Appendix Section E.
For trained models (BERT, DistilBERT, and GPT-2), a higher proportion of neurons (up to 50%) can be ablated with a relatively minor impact on primary task performance and minimal interference with auxiliary concepts. This suggests substantial neuronal redundancy, wherein multiple neurons appear to encode overlapping features.
Datasets. We consider various classification based tasks; sentiment analysis (IMDB, (Maas et al., 2011)), (SST2, (Socher et al., 2013)), emotion detection (Dair-Ai/Emotions Saravia et al. (2018)), news classification (AG-News (Zhang et al., 2015)) and article content categorization (DBPedia-14 (Zhang et al., 2015)).
Table 3: Evaluation of selected models on IMDB, SST2, AG-News, and DBPedia-14 datasets using activation range and neuron masking techniques. Performance metrics are calculated using class level 10% trimmed means at the class level. Metrics are detailed in Section 2.1. For GPT-2 and Bert 50% and for Llama-3.2-3B 30% neurons selected.
Model Dataset Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
BERT IMDB 0.928 0.904 0.928 0.904 -0.190 -0.353 0.059 -0.078 -0.184 -0.360 0.058 0.030
SST2 0.910 0.903 0.910 0.903 -0.051 -0.313 0.031 -0.046 -0.060 -0.330 0.031 0.043
AG-NEWS 0.948 0.929 0.948 0.929 -0.271 -0.590 0.012 -0.074 -0.261 -0.590 0.013 -0.009
Emotions 0.894 0.834 0.917 0.876 -0.291 -0.633 0.013 -0.265 -0.279 -0.635 0.014 -0.069
DBPedia-14 0.992 0.991 0.990 0.989 -0.028 -0.786 $0$0.017 -0.015 -0.766 $0$0.000
GPT-2 IMDB 0.952 0.939 0.952 0.939 -0.196 -0.188 0.033 0.045 -0.195 -0.197 0.031 0.042
SST2 0.966 0.958 0.966 0.958 -0.165 -0.190 0.025 0.032 -0.159 -0.192 0.025 0.028
AG-NEWS 0.945 0.933 0.945 0.933 -0.871 -0.877 -0.155 -0.163 -0.849 -0.862 -0.063 -0.223
Emotions 0.905 0.892 0.930 0.919 -0.735 -0.738 -0.103 -0.103 -0.737 -0.739 -0.044 -0.046
DBPedia-14 0.993 0.990 0.990 0.988 -0.810 -0.845 -0.154 -0.177 -0.782 -0.825 -0.015 -0.031
Llama IMDB 0.952 0.939 0.952 0.939 -0.196 -0.188 0.033 0.045 -0.195 -0.197 0.031 0.042
SST2 1.000 0.559 1.000 0.559 -0.760 -0.429 -0.394 -0.295 -0.756 -0.427 -0.384 -0.291
AG-NEWS 1.000 0.744 1.000 0.744 -0.934 -0.725 -0.660 -0.572 -0.935 -0.725 -0.484 -0.454
Emotions 0.815 0.472 0.823 0.477 -0.795 -0.470 -0.696 -0.429 -0.797 -0.469 -0.594 -0.404
DBPedia-14 1.000 0.533 1.000 0.563 -0.992 -0.528 -0.912 -0.445 -0.986 -0.527 -0.663 -0.354
5.2 Results and Analysis
Table 3 presents results for the concept removal task across five benchmark datasets (Class-wise detailed results are provided in Appendix Section J), demonstrating the effectiveness of our range-based masking approach compared to traditional neuron masking.
On binary classification tasks (IMDB, SST2), both masking approaches show moderate performance drops in targeted concepts. This suggests higher redundancy for coarser binary concepts. Multi-class classification tasks with fine-grained labels, such as AG-News, Emotions, and DBPedia-14, exhibit more pronounced effects under intervention. Range-based masking results in significant degradation of primary task performance while preserving auxiliary concept accuracy, this is particularly evident in results for AG-News.
GPT-2, despite being fine-tuned but trained in an autoregressive manner, shows substantially higher vulnerability with major drops in AG-NEWS ($\Delta_{acc}=-0.849$) and DBPedia-14 ($\Delta_{acc}=-0.782$). This increased sensitivity may be attributed to its autoregressive training objective, which potentially leads to more sequential and less redundant concept encodings. The Llama-3.2-3B model, evaluated in a few-shot setting without task-specific training, experiences the most severe degradation across all datasets (often exceeding $-0.90$), suggesting that pre-trained representations without task-specific fine-tuning are more vulnerable to targeted neuron interventions.
Table 4 in Appendix Section C presents the results showing the impact of concept erasure intervention on latent LLM capabilities such as fluency and generalization. Neuron masking degrades performance, increasing perplexity by (3.8-5.74) and lowering MMLU accuracy. In contrast, activation range masking raises perplexity by (0.5-1.1) points only, while preserving or improving MMLU scores indicates more precise and less disruptive removal.
Alternative activation interventions, beyond zeroing out, are explored in Appendix Section D, including the dampening method (Suau et al., 2024) and mean replacement (Suau et al., 2021). While these methods aim to manipulate without moving too far from the original representation, they exhibit limitations when applied to neurons. Specifically, neuron dampening increases perplexity by 2.9–3.7 points and often degrades MMLU accuracy (up to -0.045), whereas range-based dampening confines perplexity increases to 0.5–0.8 points and occasionally improves MMLU (up to +0.035). Similarly, mean replacement leads to substantial degradation when applied to neurons (perplexity increases of 7.4–8.8), while range-restricted replacement reduces the impact to below 0.7 points.
However, all approaches suffer from rigid static suppression or substitution, which fail to account for concept-specific activation dynamics. To address this issue, we introduce a novel adaptive dampening technique. This method modulates suppression in proportion to each activation’s deviation from its class-conditional mean, enabling data driven suppression. Adaptive dampening achieves the strongest balance across all metrics: perplexity remains low (0.41–0.61), MMLU is maintained or improved (up to +0.03), and collateral damage to auxiliary concepts is minimized (CAcc drops consistently below -0.3, often under -0.15), outperforming dampening, mean replacement and zeroing out approaches.
These results demonstrate that precise intervention in specific activation ranges, enables significantly more targeted concept manipulation while preserving auxiliary concepts, highlighting how conceptual information is encoded within specific activation patterns rather than isolated to individual neurons and underscoring the importance of activation ranges in capturing neuron-concept relationships.
Figure 6: Accuracy comparisons between Neuronal Range manipulation (green) and complete neuron manipulation (orange) methods on GPT-2 model.
Percentage Masking Effect
As more neurons are masked, performance gains of range-based masking over the neuron masking baseline become increasingly evident. Beyond a critical threshold of the number of masked neurons, baseline performance degrades sharply, while our method remains stable up to masking of 100% neurons. This arises from two factors: (1) models have a large number of polysemantic neurons and higher masking rates increase the chance of ablating them, and (2) blocking/manipulating a higher percentage of the model’s neurons creates a significant deviation from the original model’s behavior. For low-activation neurons with respect to the concept of interest, discrete neuron masking i.e. completely masking out a neuron, becomes unreliable, as shown in Figure 6, with a steep performance drop after masking 50% neurons. This underscores the need for finer-grained attribution; our range-based method offers such precision, preserving model behaviour under extensive masking.
The relatively stable results on auxiliary concept when using range-based masking at high percentage of neurons reduces the need to find an optimum threshold for the number of neurons to ablate which is critical to neuron masking.
6 Related Work
While we have discussed closely related approaches in Section 2, here, we briefly review additional relevant techniques. Circuit discovery identifies groups of neurons that jointly encode concepts, providing a structured view of model behavior (Marks et al., 2024; Conmy et al., 2023; Olah et al., 2020). However, extracting circuits is computationally intensive and lacks fine-grained neuron-level attribution. Gradient-based methods attribute predictions to input features by tracking gradients through the network, with integrated gradients (Sundararajan et al., 2017; Dai et al., 2022) being a widely used approach. However, they struggle with polysemanticity, as they do not disentangle overlapping concepts within neurons. Causal analysis methods intervene on internal components to assess their role in encoding concepts. Causal tracing measures the effect of corrupting activations on model performance (Vig et al., 2020; Meng et al., 2022), while causal mediation analysis quantifies information propagation through neurons (Vig et al., 2020). Although effective, these techniques require costly perturbation experiments. Beyond neuron-level analysis, representation-level methods examine hidden states and their relationship to model outputs and concepts (Veldhoen et al., 2016; Tenney et al., 2019; Liu et al., 2019). Sparse probing (Gurnee et al., 2023) compresses hidden representations into sparse, interpretable subspaces. While prior work has advanced interpretability, most methods rely on discrete neuron-to-concept mappings, which fail to account for polysemanticity (Sajjad et al., 2022). Our work extends activation-based approaches by introducing activation ranges as the unit of interpretability to enable more precise concept attribution and intervention.
7 Conclusion
In this work, we challenged traditional assumptions about neuron interpretability by reframing polysemanticity as a resource rather than a limitation in interpreting neurons. Through an in depth analysis, we uncovered that neuronal activations for individual concepts exhibit distinct, Gaussian-like distributions. This discovery allows for a more precise understanding of how neurons encode multiple concepts, enabling us to move beyond coarse, monolithic neuron-to-concept mappings. Building upon these insights, we proposed $\mathsf{NeuronLens}$, a novel range-based framework for neuron interpretation and manipulation. $\mathsf{NeuronLens}$ offers fine-grained control that reduces interference with unrelated concepts by attributing specific activation ranges within neurons to individual concepts. Extensive empirical evaluations demonstrated that $\mathsf{NeuronLens}$ outperforms neuronal attribution methods in maintaining concept-specific precision while minimizing unintended side effects. Notably, while targeted concept removal remains equally effective when comparing neuron vs range based interventions, our approach achieves superior preservation of auxiliary concepts without compromising the primary goal. An important direction for future work is exploring our range-based method as a metric for quantifying polysemanticity in neural networks. This approach may also serve as a diagnostic tool to evaluate the effectiveness of sparse autoencoders (SAEs) in disentangling concept representations across individual neurons.
Appendix A Impact Statement
This work advances neural network interpretability by providing a fine-grained understanding of concept encoding in language models. The proposed $\mathsf{NeuronLens}$ framework enables precise control of model behavior, benefiting research in model safety and reliability. While this improved understanding could potentially be misused, the work’s theoretical nature and focus on interpretability methods makes immediate harmful applications unlikely.
Appendix B Limitations
While $\mathsf{NeuronLens}$ can disentangle polysemanticity to a degree using Gaussian Like Distribution, it is unable to completely disentangle concepts encoded in the polysemantic neurons, because there still is a significant overlap in the distributions of concepts in activations. Additionally, in this work, we use $\tau$ to be a fixed value of 2.5 to make the comparison of approaches fair, but $\tau$ selection can be optimized to be more sophisticated. We also get results primarily from the penultimate layer, and not the intermediate or earlier layers, however, we do give ablation and rationale for this choice in Appendix I
Appendix C General LLM Capabilities
We evaluate the general capabilities of large language models (LLMs) using the MMLU benchmark [Hendrycks et al., 2021] and perplexity scores on Wikipedia texts [Foundation, ]. Table 4 presents the comparative performance of neuron masking and activation range masking. Neuron masking leads to notable increases in perplexity, exceeding 3 points in the best case, whereas range masking results in a maximum increase of only 1.1. In terms of MMLU accuracy, neuron masking consistently reduces performance across all settings, while range masking preserves or improves performance in most cases, with degradation observed in only one instance.
Table 4: Evaluation of LLMs latent capabilities using Wikipedia for perplexity and zeroshot MMLU for testing generalisation capabilities. Llama-3.2-3B model
Dataset Base Values Neuron Masking Activation Range Masking
Perplexity MMLU Perplexity MMLU Perplexity MMLU
IMDB 7.007 0.530 10.990 0.515 7.550 0.530
SST2 7.007 0.530 11.688 0.510 8.150 0.537
AG-NEWS 7.007 0.530 12.757 0.510 8.022 0.533
Emotions 7.007 0.530 11.630 0.526 8.063 0.526
DBPedia-14 7.007 0.530 12.230 0.507 7.903 0.535
Appendix D Activation Intervention
In the main text, we primarily presented results using a “zeroing out” strategy for neuron manipulation. This approach was chosen to compare neuron manipulation against range-based manipulation. However, zeroing out is considered a suboptimal strategy [Suau et al., 2024]. The primary concern with standard zeroing-out approaches is that they distort the activation distribution significantly, diverging from that of the original model. However, our range-based method selectively zeroes out only a narrow slice of the activation spectrum, thereby mitigating the adverse effects associated with hard erasure.
In this section, we explore alternative, more optimized strategies for concept removal. We also introduce a novel range-based scaling strategy that has demonstrated superior results.
Below, we explore various activation intervention strategies, comparing traditional neuron-level approaches with the nuanced range-based technique. Our comprehensive evaluation reveals that range-based manipulations consistently outperform neuron interventions across multiple metrics, with significantly less disruption to the model’s general capabilities.
Among all techniques examined, our novel adaptive dampening approach emerges as the most effective, maintaining targeted concept suppression while minimizing collateral impact on auxiliary concepts and preserving overall language modelling capabilities. This pattern holds true across different intervention methods including zeroing out, dampening, and mean replacement strategies.
D.1 Dampening
In their work, Suau et al. [2024] propose using a dampening function rather than setting neuron activations to zero outright. This approach, referred to as DAMP, corresponds to a specific choice of the intervention function $\phi(x)=\alpha x$, where $0\leq\alpha\leq 1$. In this formulation, the activations of selected neurons are scaled down by a factor $\alpha$ instead of being completely suppressed. Here, $x$ represents neuron activation. The rationale behind dampening is that a fixed intervention (like zeroing out) can disrupt the LLM’s inference dynamics, especially when a large number of neurons ($k$) are involved, thereby limiting its effectiveness. Dampening offers a less destructive intervention by allowing contextual signals to continue passing through the network. This, in turn, permits intervention on a larger set of expert neurons, potentially achieving stronger mitigation of the targeted concept.
Table 5: Evaluation of Llama-3.2-3B on a DBPedia-14 dataset using neuron and range masking techniques. 30% neurons were selected. Dampening factor used is $a=0.125$. Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Neuron Masking and Activation Range Masking show deviations from the baseline performance. PPL $\Delta$ and MMLU $\Delta$ show changes in perplexity and MMLU scores, respectively.
Class Base Values Neuron Masking (Deviations) Activation Range Masking (Deviations)
Acc Conf CAcc CConf Acc Conf CAcc CConf PPL $\Delta$ MMLU $\Delta$ Acc Conf CAcc CConf PPL $\Delta$ MMLU $\Delta$
Class 0 1.000 0.576 1.000 0.563 -0.919 -0.545 -0.281 -0.309 3.161 -0.020 -0.924 -0.545 -0.276 -0.285 0.640 -0.010
Class 1 1.000 0.526 1.000 0.567 -0.988 -0.467 -0.246 -0.270 3.578 -0.015 -0.805 -0.466 -0.193 -0.206 0.725 0.015
Class 2 1.000 0.441 1.000 0.575 -0.864 -0.391 -0.461 -0.323 2.891 -0.030 -0.869 -0.389 -0.346 -0.282 0.718 0.005
Class 3 1.000 0.461 1.000 0.573 -0.974 -0.439 -0.411 -0.346 3.036 -0.025 -0.970 -0.438 -0.282 -0.283 0.653 0.010
Class 4 1.000 0.839 1.000 0.541 -0.382 -0.597 -0.367 -0.317 2.997 $0$0.382 -0.597 -0.334 -0.284 0.691 0.020
Class 5 1.000 0.339 1.000 0.568 -0.970 -0.326 -0.239 -0.246 3.503 0.010 -0.970 -0.325 -0.197 -0.187 0.810 0.015
Class 6 1.000 0.810 1.000 0.545 -0.233 -0.638 -0.194 -0.276 3.126 -0.010 -0.241 -0.637 -0.174 -0.203 0.697 -0.010
Class 7 1.000 0.595 1.000 0.562 -0.210 -0.382 -0.206 -0.226 3.037 $0$0.179 -0.376 -0.123 -0.143 0.546 0.015
Class 8 1.000 0.417 1.000 0.574 -0.310 -0.416 -0.335 -0.297 3.001 0.020 -0.346 -0.416 -0.200 -0.187 0.624 0.015
Class 9 1.000 0.526 1.000 0.567 -0.820 -0.465 -0.327 -0.264 3.369 -0.030 -0.809 -0.463 -0.213 -0.189 0.596 0.000
Class 10 1.000 0.505 1.000 0.569 -0.691 -0.466 -0.389 -0.314 3.732 $0$0.696 -0.465 -0.267 -0.198 0.695 -0.015
Class 11 1.000 0.497 1.000 0.569 -0.873 -0.432 -0.472 -0.289 3.070 -0.030 -0.865 -0.427 -0.335 -0.205 0.594 -0.015
Class 12 1.000 0.573 1.000 0.563 -0.720 -0.452 -0.295 -0.221 3.410 -0.045 -0.723 -0.451 -0.190 -0.163 0.595 0.035
Class 13 1.000 0.567 1.000 0.564 -0.951 -0.537 -0.226 -0.189 2.995 $0$0.955 -0.536 -0.157 -0.150 0.672 0.005
Table 5 presents a comparative analysis of two intervention strategies, neuron masking and activation range masking, when employing the Dampening technique with $\alpha=0.5$. The evaluation spans 14 classes and utilizes the metrics: accuracy (Acc), confidence (Conf), class-wise accuracy (CAcc), class-wise confidence (CConf), alterations in perplexity (PPL), and MMLU score.
A consistent trend emerges across the primary metrics (Acc, Conf, CAcc, and CConf), where activation range masking demonstrates superior performance over neuron masking. Interventions based on activation ranges lead to a notably smaller decline in the accuracy and confidence associated with auxiliary concepts. For example, in Class 3, while neuron masking results in an accuracy drop of -0.974 in the targeted class and auxiliary class accuracy decrease of -0.411, activation range masking, despite a comparable accuracy reduction in the targeted class (-0.970), shows a less severe impact on auxiliary class accuracy (-0.283). This pattern of activation range masking better preserves auxiliary class performance, is evident across all evaluated classes.
Examining the broader effects on language modeling capabilities reveals significant distinctions between the two approaches. Neuron masking results in a considerable rise in perplexity (PPL), with increases ranging from +2.891 to +3.732 across all the classes. Furthermore, it tends to cause more pronounced negative shifts in MMLU scores, reaching as low as -0.045. Conversely, activation range masking results in substantially smaller increments in perplexity, falling within the +0.546 to +0.810 range, and frequently results in improved or minimally altered MMLU scores, with gains up to +0.035.
D.2 Mean Replacement
Another approach of activation replacement discussed in the literature[Suau et al., 2021] is replacing it with the mean activation value. We provide the results for this type of replacement in Table 6.
The mean replacement strategy corresponds to setting the intervention function to $\phi(x)=\mu$, where $\mu$ is the mean activation of the neuron $x$ computed over a general next-token prediction task on the Wikipedia[Foundation, ].
Table 6: Evaluation of Llama-3.2-3B on a DBPedia-14 dataset using neuron and range masking techniques. 30% neurons were selected. Mean Activation $\mu$ is used as replacement value. Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Neuron Masking and Activation Range Masking show deviations from the baseline performance. PPL $\Delta$ and MMLU $\Delta$ show changes in perplexity and MMLU scores, respectively.
Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf PPL $\Delta$ MMLU $\Delta$ Acc Conf CAcc CConf PPL $\Delta$ MMLU $\Delta$
Class 0 1.000 0.576 1.000 0.563 -1.000 -0.576 -0.685 -0.554 7.681 -0.025 -1.000 -0.576 -0.551 -0.545 0.687 -0.005
Class 1 1.000 0.526 1.000 0.567 -1.000 -0.526 -0.554 -0.550 8.437 -0.030 -1.000 -0.526 -0.356 -0.517 0.583 0.015
Class 2 1.000 0.441 1.000 0.575 -0.995 -0.441 -0.697 -0.556 7.567 -0.015 -0.995 -0.440 -0.574 -0.536 0.520 -0.010
Class 3 1.000 0.461 1.000 0.573 -1.000 -0.461 -0.766 -0.561 8.005 -0.015 -1.000 -0.461 -0.538 -0.534 0.543 0.010
Class 4 1.000 0.839 1.000 0.541 -1.000 -0.838 -0.724 -0.528 8.239 0.010 -0.995 -0.838 -0.502 -0.503 0.565 0.005
Class 5 1.000 0.339 1.000 0.568 -1.000 -0.339 -0.616 -0.551 7.753 0.010 -1.000 -0.339 -0.382 -0.510 0.552 0.005
Class 6 1.000 0.810 1.000 0.545 -0.313 -0.805 -0.549 -0.531 7.880 -0.005 -0.292 -0.805 -0.336 -0.499 0.547 0.020
Class 7 1.000 0.595 1.000 0.562 -1.000 -0.592 -0.491 -0.535 7.413 -0.010 -0.995 -0.591 -0.267 -0.449 0.462 0.000
Class 8 1.000 0.417 1.000 0.574 -0.928 -0.414 -0.632 -0.556 7.688 0.015 -0.934 -0.414 -0.298 -0.489 0.495 0.015
Class 9 1.000 0.526 1.000 0.567 -1.000 -0.526 -0.611 -0.544 8.057 -0.035 -1.000 -0.526 -0.370 -0.482 0.467 0.015
Class 10 1.000 0.505 1.000 0.569 -0.998 -0.505 -0.642 -0.558 8.791 -0.020 -0.998 -0.505 -0.406 -0.485 0.484 0.005
Class 11 1.000 0.497 1.000 0.569 -1.000 -0.497 -0.719 -0.543 7.903 0.025 -1.000 -0.497 -0.447 -0.459 0.397 -0.005
Class 12 1.000 0.573 1.000 0.563 -0.904 -0.572 -0.629 -0.543 8.046 -0.005 -0.896 -0.571 -0.375 -0.484 0.425 0.000
Class 13 1.000 0.567 1.000 0.564 -1.000 -0.566 -0.526 -0.533 7.543 -0.025 -0.998 -0.566 -0.341 -0.481 0.464 -0.010
In Table 6, we assess the effect of mean replacement using both neuron masking and activation range masking. In every class, neuron masking results in more severe degradation than range-based masking across all auxiliary and general metrics.
Across metrics (Acc, Conf, CAcc, and CConf), activation range masking consistently outperforms neuron masking. The degradation in accuracy and confidence of auxiliary concepts is significantly lower under range-based interventions. For instance, in Class 3, neuron masking causes a drop of -1.000 in Acc and -0.766 in CAcc, whereas activation range masking yields a similar Acc drop (-1.000) but a substantially smaller decline in CAcc (-0.538). A similar pattern repeats across all classes; for example, in Class 0, neuron masking results in CAcc of -0.685 while activation range masking yields -0.551. In Class 7, neuron masking shows a CAcc of -0.491 compared to -0.267 for activation range masking.
Beyond auxiliary class performance, we observe substantial differences in how the two masking methods affect general language modelling capabilities. Neuron masking leads to a large increase in perplexity (PPL), ranging from +7.413 to +8.791 which is catastrophic, across classes, and induces more negative shifts in MMLU scores (as low as -0.035 for Class 9, and also for Class 0 with -0.025, Class 1 with -0.030, Class 10 with -0.020, and Class 13 with -0.025). In contrast, activation range masking results in substantially smaller increases in perplexity (+0.397 to +0.687) and often yields improved or near-zero changes in MMLU scores (up to +0.020 for Class 6, and several positive values like +0.015 for Class 1, Class 8, and Class 9).
D.3 Adaptive Dampening
We propose a novel replacement method in which the intervention function $\phi(x)$ applies runtime-controlled dampening based on the distance of the observed activation $x$ from the center of a predefined activation range. Specifically, the dampening factor $a(x)$ is linearly scaled according to the distance of $x$ from the mean $\mu$ of the neuron’s activation distribution, within the range $[\mu-2.5\sigma,\mu+2.5\sigma]$.
Let $\beta\in[0,1]$ denote the maximum dampening factor applied at the range boundaries. Then:
$a(x)=\beta\cdot\frac{|x-\mu|}{2.5\sigma},\quad\text{and}\quad\phi(x)=a(x)\cdot x.$
This ensures that when $x=\mu$ (the center of the range), $a(x)=0$ and the activation is fully suppressed via $\phi(x)=0$. At the boundaries ($x=\mu\pm 2.5\sigma$), $a(x)=\beta$, and the activation is minimally dampened. Values within the range are scaled proportionally based on their normalized distance from the mean. This adaptive dampening mechanism suppresses values near the mean while preserving those closer to the range edges.
The dampening factor $\beta$ can be optimized for different neurons based on the concept information that neuron provides. For this work, we use $\beta=0.5$ across all neurons.
Table 7: Evaluation of Llama-3.2-3B on a DBPedia-14 dataset using neuron and range masking techniques. 30% neurons were selected. Adaptive Dampening factor used is $\beta=0.5$. Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Neuron Masking and Activation Range Masking show deviations from the baseline performance. PPL $\Delta$ and MMLU $\Delta$ show changes in perplexity and MMLU scores, respectively.
Class Base Values Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf PPL $\Delta$ MMLU $\Delta$
Class 0 1.000 0.576 1.000 0.563 -0.927 -0.543 -0.215 -0.217 0.487 -0.015
Class 1 1.000 0.526 1.000 0.567 -0.791 -0.451 -0.134 -0.109 0.543 0.000
Class 2 1.000 0.441 1.000 0.575 -0.828 -0.380 -0.277 -0.215 0.540 -0.010
Class 3 1.000 0.461 1.000 0.573 -0.958 -0.432 -0.230 -0.214 0.492 0.010
Class 4 1.000 0.839 1.000 0.541 -0.346 -0.579 -0.261 -0.218 0.521 0.015
Class 5 1.000 0.339 1.000 0.568 -0.960 -0.319 -0.140 -0.116 0.609 -0.015
Class 6 1.000 0.810 1.000 0.545 -0.236 -0.613 -0.130 -0.122 0.524 -0.010
Class 7 1.000 0.595 1.000 0.562 -0.243 -0.388 -0.108 -0.080 0.408 0.005
Class 8 1.000 0.417 1.000 0.574 -0.440 -0.414 -0.152 -0.088 0.465 0.030
Class 9 1.000 0.526 1.000 0.567 -0.799 -0.459 -0.182 -0.131 0.445 0.005
Class 10 1.000 0.505 1.000 0.569 -0.684 -0.451 -0.222 -0.130 0.513 -0.010
Class 11 1.000 0.497 1.000 0.569 -0.836 -0.420 -0.308 -0.155 0.440 -0.005
Class 12 1.000 0.573 1.000 0.563 -0.720 -0.451 -0.172 -0.095 0.444 0.025
Class 13 1.000 0.567 1.000 0.564 -0.941 -0.530 -0.142 -0.098 0.502 0.010
In Table 7 we evaluate the adaptive dampening variant of the replacement function. This approach outperforms both neuron masking and static activation masking across all metrics.
In auxiliary class metrics, adaptive dampening yields much smaller degradation. Auxilary class accuracy (CAcc) and confidence (CConf) show significantly reduced drops compared to other methods. For example, in Class 0, CAcc drops only $-0.215$ compared to $-0.685$ under neuron masking and $-0.551$ under hard activation masking. The effect is consistent across classes, with most CAcc and CConf drops staying well below $-0.3$, and in many cases below $-0.15$.
Language modeling metrics show this approach to be exceptionally efficient. Perplexity increases are minimal, remaining within $+0.408$ to $+0.609$, substantially lower than all hard-masking variants. MMLU deltas also stay close to zero, with several classes showing improvement (e.g., Class 8: $+0.030$, Class 4: $+0.015$). Notably, no class suffers significant MMLU degradation.
Appendix E Training Details
For BERT, DistilBERT, and Llama, we utilize pretrained models. Since BERT, and DistilBert are not inherently trained as a conversational agent, we use top-performing fine-tuned models from the Hugging Face repository. For the Llama model, few-shot prompt completion is employed to predict class labels. This involves providing a small number of training samples from the dataset to guide the model’s predictions.
For GPT-2, we fine-tune the pretrained model across all datasets for three epochs. The input sequence is constructed by concatenating the text with a <sep> token, followed by the class label, and ending with an <eos> token. During training, the loss is back-propagated only for the class label token, while all other tokens are assigned a skip label (-100). Additionally, all class labels are added to the model’s dictionary as special single-token entries.
In the case of Bert-based models, record the activation of the CLS token, In the case of GPT-2 and Llama models, we record the last token output when the class token is being predicted. The intervention is applied to the appropriate token on the residual stream.
Dataset Preprocessing for Llama For Llama we process whole datasets in few shout settings and only curate 2000 samples per class, where the model prediction was correct.
Appendix F Compute Details
All experiments, including activation extraction and interventions on large language models (LLMs), were conducted using an NVIDIA RTX 3090 GPU equipped with 24GB of VRAM. 64GB RAM.
Appendix G Saliency details
Max activations. Frankle and Carbin [2019] extract high neural activations as a saliency ranking metric relying upon the rationale that maximally activating neurons are salient as these neurons play a critical role in controlling the model’s output, highlighting their importance for a concept $c$.To identify them, the column-wise mean of absolute neuronal activations in $H^{l}_{c}$, $H^{l}_{c}$ is defined in Section2.3, is computed, given that high negative activations also carry significant signals [Voita et al., 2023]. The magnitude of the means is then considered as a ranking for concept $c$.
Probe analysis. Dalvi et al. [2019b] train a linear classifier on the hidden representations $H^{l}_{c}$ to distinguish between concepts. The learned model weights are then utilized as a saliency ranking. This process involves learning a weight matrix $W\in\mathbb{R}^{d\times|c|}$, where $d$ is the hidden dimension and $|c|$ is the number of concept classes. The absolute weight values of each row in the weight matrix are used as a ranking for the importance of each neuron for a given concept. To prevent the emergence of redundant solutions characterized by minimal variations in the weights, the probe is trained using the elastic regularization technique.
Probeless. Antverg and Belinkov [2022] examine individual neurons, without the need for auxiliary classifiers, using the element-wise difference between mean vectors. The element-wise difference between mean vectors is computed as $r=\sum_{c,c^{\prime}\in C}|q(c)-q(c^{\prime})|$, where $r\in\mathbb{R}^{d}$ and $d$ is the hidden dimension. The final neuron saliency ranking is obtained by sorting $r$ in descending order.
Table 8: Performance drops relative to Baseline configuration (i.e.: unaltered model’s performance) for three techniques: Probeless, Probe, and Max. All values show the difference from Base values. Results are for Emotions dataset on the GPT-2 model using 30% salient neurons of each method. Metrics are detailed in 2.1.
Probeless Probe Max
Class Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
Class 0 -0.738 -0.733 -0.103 -0.097 -0.613 -0.650 -0.010 -0.038 -0.695 -0.751 -0.125 -0.124
Class 1 0.045 0.041 -0.113 -0.112 -0.014 -0.015 -0.010 -0.034 -0.879 -0.882 -0.019 -0.009
Class 2 -0.570 -0.541 -0.052 -0.057 0.017 0.009 -0.347 -0.359 -0.776 -0.736 -0.029 -0.032
Class 3 -0.164 -0.166 -0.035 -0.038 0.078 0.061 -0.047 -0.104 -0.713 -0.714 -0.006 -0.007
Class 4 -0.623 -0.617 -0.087 -0.084 -0.005 -0.010 -0.003 -0.020 -0.754 -0.753 -0.240 -0.248
Class 5 -0.817 -0.714 -0.101 -0.105 -0.206 -0.127 0.003 -0.010 -0.587 -0.601 -0.301 -0.308
Appendix H Hyperparameter Ablation
For target concept, $\tau$ values $0.3-2.4$ show decreasing accuracy/confidence, stabilizing at $\tau=2.4$ (accuracy $0.6126$). Beyond $2.4$, negligible additional degradation occurs, indicating we’ve captured the complete target concept activation range. Importantly, while target performance stabilizes after $\tau=2.4$, auxiliary task performance declines after $\tau=2.7$. Complement accuracy stays above 0.93 until then before dropping to 0.8795 at $\tau=4.5$. This aligns with normal distribution properties where 95-99% of values fall within ±2.5 standard deviations.
Table 9: Performance metrics for varying $\tau$ values.
$\tau$ Acc Conf CAcc CConf
0.3 0.9021 0.8858 0.9452 0.9358
0.6 0.8439 0.8185 0.9424 0.9327
0.9 0.7801 0.7486 0.9391 0.9263
1.2 0.7295 0.6950 0.9340 0.9174
1.5 0.6834 0.6482 0.9337 0.9093
1.8 0.6424 0.6141 0.9331 0.9000
2.1 0.6184 0.5926 0.9327 0.8910
2.4 0.6126 0.5858 0.9314 0.8846
2.7 0.6024 0.5798 0.9280 0.8800
3.0 0.5971 0.5776 0.9234 0.8777
3.3 0.5963 0.5786 0.9173 0.8753
3.6 0.5970 0.5794 0.9097 0.8729
3.9 0.5976 0.5802 0.9020 0.8698
4.2 0.5967 0.5798 0.8908 0.8642
4.5 0.5967 0.5798 0.8795 0.8577
Appendix I Layer Ablation
I.1 Statistical Results
We analyze concept level activation distributions across all 12 layers of GPT-2, measuring kurtosis (where a value of 3.0 indicates a Gaussian distribution), skewness (where 0 indicates symmetry), and practical normality in Table 10:
Table 10: Statistical analysis of different layers showing skewness, kurtosis, and Kolmogorov- Smirnov test results. GPT2 model. AG-News Dataset
Layer Kurtosis Skewness Practical Normality(10%)
1 3.9314 0.0430 0.7913
2 3.7622 -0.0091 0.9525
3 3.4109 -0.0143 0.9870
4 3.5582 -0.0073 0.9801
5 3.6145 0.0051 0.9730
6 3.5318 0.0086 0.9769
7 3.3461 0.0083 0.9880
8 3.2763 0.0037 0.9870
9 3.2267 0.0039 0.9860
10 3.2057 0.0029 0.9899
11 3.2105 -0.0002 0.9912
12 3.2061 -0.0014 0.9919
These results show that kurtosis values converge toward 3.0 (the Gaussian ideal) as layers progress, skewness values remain near zero across all layers, and practical normality scores are close to 1 across all layers. Importantly, if the activations were not clustered into continuous intervals and were in disconnected islands of activations, these would be reflected in the score for the practical normality and other statistical metrics.
I.2 Qualitative Results
We expanded our visualization approach shown in Figures figs. 8, 8, 10, 10, 12, 12, 14, 14, 16, 16, 18 and 18 in Figure 4) to all layers in the model. The visualizations demonstrate an interesting progression: while all layers exhibit Gaussian-like distributions on the class level, class concepts aren’t separated in the activation spectrum Gaussians of the early layers. This aligns with the understanding that lower layers capture more basic features rather than high-level semantic features like class. However, distinct concept-level Gaussian distributions begin forming as early as layers 5-6, becoming increasingly separable in deeper layers.
Figure 7: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 1
Figure 8: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 2
Figure 9: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 3
Figure 10: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 4
Figure 11: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 5
Figure 12: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 6
Figure 13: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 7
Figure 14: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 8
Figure 15: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 9
Figure 16: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 10
Figure 17: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 11
Figure 18: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 12
I.3 Masking Results
In Table 11 and Table 12 we provide results of applying both approaches on all layers of GPT-2 model on Emotions dataset. From the results we can see that: Early layers (1-3) show highly variable and often severe impacts: Layer 1 exhibits minimal effects ($\Delta Acc=-0.113$, $\Delta CAcc=-0.064$), while Layers 2-3 show extreme degradation ($\Delta Acc\approx-0.7$, $\Delta CAcc>-0.5$). Middle layers (4-8) demonstrate inconsistent behavior with high variance in impacts. Layer 12, however, achieves an optimal balance: it maintains substantial primary task impact ($\Delta Acc=-0.571$) while minimizing auxiliary concept interference ($\Delta CAcc=-0.060$). This pattern holds true for both neuron masking and range masking techniques, with range masking showing slightly better preservation of auxiliary concepts ($\Delta CAcc=-0.045$). The mid-range primary task degradation combined with minimal auxiliary impact makes Layer 12 the most suitable for targeted interventions, offering better control and specificity compared to earlier layers.
Table 11: Evaluation of layer selection on GPT-2 model on the Emotions dataset using neuron and range masking techniques. 20% Neurons selected. Here, Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance.
Layer Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
1 Class 0 0.970 0.957 0.915 0.904 -0.029 -0.071 -0.074 -0.100 0.006 0.002 -0.004 -0.005
Class 1 0.933 0.932 0.931 0.913 -0.011 -0.056 -0.090 -0.116 0.001 -0.003 -0.004 -0.004
Class 2 0.901 0.865 0.934 0.924 -0.206 -0.195 -0.052 -0.092 -0.019 -0.015 -0.001 -0.002
Class 3 0.926 0.924 0.932 0.919 -0.128 -0.152 -0.051 -0.090 -0.004 -0.005 -0.001 -0.002
Class 4 0.885 0.867 0.938 0.927 -0.055 -0.084 -0.061 -0.093 -0.016 -0.009 0.002 -0.001
Class 5 0.851 0.786 0.934 0.924 -0.249 -0.217 -0.055 -0.094 0.016 0.013 -0.004 -0.005
2 Class 0 0.970 0.957 0.915 0.904 -0.804 -0.808 -0.389 -0.386 -0.061 -0.133 -0.077 -0.096
Class 1 0.933 0.932 0.931 0.913 0.053 -0.003 -0.819 -0.781 -0.011 -0.049 -0.110 -0.145
Class 2 0.901 0.865 0.934 0.924 -0.868 -0.737 -0.515 -0.519 -0.365 -0.337 -0.077 -0.126
Class 3 0.926 0.924 0.932 0.919 -0.870 -0.805 -0.498 -0.501 -0.215 -0.248 -0.096 -0.153
Class 4 0.885 0.867 0.938 0.927 -0.729 -0.707 -0.461 -0.463 -0.042 -0.077 -0.076 -0.116
Class 5 0.851 0.786 0.934 0.924 -0.845 -0.769 -0.511 -0.508 -0.229 -0.188 -0.106 -0.163
3 Class 0 0.970 0.957 0.915 0.904 -0.896 -0.904 -0.824 -0.832 -0.647 -0.688 -0.517 -0.544
Class 1 0.933 0.932 0.931 0.913 -0.901 -0.916 -0.835 -0.832 -0.568 -0.607 -0.609 -0.630
Class 2 0.901 0.865 0.934 0.924 -0.868 -0.845 -0.838 -0.851 -0.605 -0.600 -0.589 -0.619
Class 3 0.926 0.924 0.932 0.919 -0.868 -0.896 -0.830 -0.840 -0.567 -0.605 -0.567 -0.596
Class 4 0.885 0.867 0.938 0.927 -0.800 -0.811 -0.849 -0.857 -0.502 -0.522 -0.513 -0.544
Class 5 0.851 0.786 0.934 0.924 0.022 0.081 -0.865 -0.881 -0.155 -0.124 -0.561 -0.596
4 Class 0 0.970 0.957 0.915 0.904 -0.650 -0.703 -0.698 -0.764 -0.608 -0.621 -0.499 -0.510
Class 1 0.933 0.932 0.931 0.913 -0.845 -0.884 -0.667 -0.725 -0.491 -0.519 -0.480 -0.491
Class 2 0.901 0.865 0.934 0.924 -0.858 -0.824 -0.772 -0.809 -0.488 -0.497 -0.506 -0.523
Class 3 0.926 0.924 0.932 0.919 -0.700 -0.808 -0.663 -0.739 -0.534 -0.546 -0.512 -0.528
Class 4 0.885 0.867 0.938 0.927 -0.239 -0.514 -0.754 -0.797 -0.304 -0.307 -0.452 -0.471
Class 5 0.851 0.786 0.934 0.924 -0.612 -0.463 -0.692 -0.765 -0.047 -0.038 -0.525 -0.541
5 Class 0 0.970 0.957 0.915 0.904 -0.838 -0.852 -0.492 -0.630 -0.695 -0.688 -0.554 -0.555
Class 1 0.933 0.932 0.931 0.913 -0.387 -0.563 -0.683 -0.714 -0.552 -0.564 -0.605 -0.599
Class 2 0.901 0.865 0.934 0.924 -0.702 -0.700 -0.634 -0.690 -0.472 -0.470 -0.607 -0.605
Class 3 0.926 0.924 0.932 0.919 -0.361 -0.507 -0.615 -0.692 -0.567 -0.575 -0.538 -0.539
Class 4 0.885 0.867 0.938 0.927 -0.873 -0.844 -0.525 -0.650 -0.668 -0.653 -0.594 -0.594
Class 5 0.851 0.786 0.934 0.924 -0.637 -0.573 -0.588 -0.681 -0.069 -0.022 -0.548 -0.553
6 Class 0 0.970 0.957 0.915 0.904 -0.720 -0.775 -0.829 -0.830 -0.484 -0.499 -0.318 -0.322
Class 1 0.933 0.932 0.931 0.913 -0.871 -0.887 -0.750 -0.768 -0.176 -0.195 -0.499 -0.499
Class 2 0.901 0.865 0.934 0.924 -0.895 -0.860 -0.735 -0.773 -0.680 -0.638 -0.335 -0.348
Class 3 0.926 0.924 0.932 0.919 -0.863 -0.884 -0.772 -0.793 -0.418 -0.431 -0.379 -0.381
Class 4 0.885 0.867 0.938 0.927 -0.621 -0.669 -0.743 -0.784 -0.430 -0.435 -0.247 -0.262
Class 5 0.851 0.786 0.934 0.924 -0.143 -0.086 -0.808 -0.831 -0.114 -0.070 -0.474 -0.478
Table 12: Evaluation of layer selection on GPT-2 model on the Emotions dataset using neuron and range masking techniques. 20% Neurons selected. Here, Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance.
Layer Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
7 Class 0 0.970 0.957 0.915 0.904 -0.908 -0.901 -0.752 -0.753 -0.527 -0.538 -0.492 -0.498
Class 1 0.933 0.932 0.931 0.913 -0.884 -0.895 -0.743 -0.729 -0.484 -0.509 -0.330 -0.338
Class 2 0.901 0.865 0.934 0.924 -0.866 -0.835 -0.767 -0.765 -0.451 -0.442 -0.336 -0.355
Class 3 0.926 0.924 0.932 0.919 -0.786 -0.819 -0.641 -0.666 -0.445 -0.457 -0.331 -0.346
Class 4 0.885 0.867 0.938 0.927 -0.626 -0.618 -0.810 -0.817 -0.341 -0.335 -0.521 -0.532
Class 5 0.851 0.786 0.934 0.924 0.106 0.147 -0.810 -0.811 0.102 0.107 -0.547 -0.553
8 Class 0 0.970 0.957 0.915 0.904 -0.776 -0.791 -0.209 -0.291 -0.191 -0.312 -0.082 -0.114
Class 1 0.933 0.932 0.931 0.913 -0.585 -0.667 -0.412 -0.441 -0.591 -0.644 -0.199 -0.227
Class 2 0.901 0.865 0.934 0.924 -0.692 -0.716 -0.469 -0.496 -0.560 -0.562 -0.468 -0.486
Class 3 0.926 0.924 0.932 0.919 -0.657 -0.714 -0.415 -0.464 -0.468 -0.503 -0.230 -0.266
Class 4 0.885 0.867 0.938 0.927 -0.501 -0.509 -0.531 -0.569 -0.201 -0.234 -0.258 -0.290
Class 5 0.851 0.786 0.934 0.924 -0.092 -0.050 -0.634 -0.647 0.065 0.058 -0.279 -0.308
9 Class 0 0.970 0.957 0.915 0.904 -0.759 -0.768 -0.311 -0.351 -0.610 -0.661 -0.307 -0.328
Class 1 0.933 0.932 0.931 0.913 -0.570 -0.713 -0.319 -0.346 -0.906 -0.910 -0.267 -0.298
Class 2 0.901 0.865 0.934 0.924 -0.424 -0.520 -0.504 -0.531 -0.635 -0.643 -0.579 -0.595
Class 3 0.926 0.924 0.932 0.919 -0.810 -0.834 -0.501 -0.502 -0.759 -0.772 -0.502 -0.516
Class 4 0.885 0.867 0.938 0.927 -0.358 -0.357 -0.476 -0.481 -0.587 -0.566 -0.519 -0.527
Class 5 0.851 0.786 0.934 0.924 -0.133 -0.101 -0.546 -0.554 0.106 0.104 -0.450 -0.462
10 Class 0 0.970 0.957 0.915 0.904 -0.733 -0.741 -0.105 -0.126 -0.624 -0.659 -0.146 -0.163
Class 1 0.933 0.932 0.931 0.913 -0.389 -0.671 -0.178 -0.209 -0.899 -0.911 -0.254 -0.285
Class 2 0.901 0.865 0.934 0.924 -0.230 -0.513 -0.116 -0.224 -0.699 -0.735 -0.409 -0.451
Class 3 0.926 0.924 0.932 0.919 -0.434 -0.687 -0.081 -0.133 -0.898 -0.905 -0.401 -0.455
Class 4 0.885 0.867 0.938 0.927 -0.489 -0.506 -0.188 -0.256 -0.140 -0.186 -0.063 -0.102
Class 5 0.851 0.786 0.934 0.924 -0.306 -0.243 -0.157 -0.240 0.063 0.010 -0.095 -0.127
11 Class 0 0.970 0.957 0.915 0.904 -0.358 -0.496 -0.382 -0.414 -0.301 -0.441 -0.121 -0.148
Class 1 0.933 0.932 0.931 0.913 -0.800 -0.857 -0.078 -0.123 -0.858 -0.875 -0.128 -0.162
Class 2 0.901 0.865 0.934 0.924 -0.897 -0.861 -0.416 -0.450 -0.901 -0.864 -0.464 -0.500
Class 3 0.926 0.924 0.932 0.919 -0.923 -0.921 -0.427 -0.470 -0.913 -0.914 -0.354 -0.393
Class 4 0.885 0.867 0.938 0.927 -0.152 -0.212 -0.039 -0.075 -0.210 -0.239 -0.181 -0.204
Class 5 0.851 0.786 0.934 0.924 0.047 -0.028 -0.131 -0.173 0.053 0.002 -0.142 -0.159
12 Class 0 0.970 0.957 0.915 0.904 -0.550 -0.603 -0.013 -0.003 -0.542 -0.594 0.005 0.012
Class 1 0.933 0.932 0.931 0.913 -0.526 -0.545 0.001 0.012 -0.521 -0.538 -0.005 -0.004
Class 2 0.901 0.865 0.934 0.924 -0.416 -0.402 0.002 0.006 -0.419 -0.407 0.007 0.006
Class 3 0.926 0.924 0.932 0.919 -0.561 -0.576 -0.007 0.003 -0.561 -0.572 0.000 0.005
Class 4 0.885 0.867 0.938 0.927 -0.655 -0.658 -0.042 -0.034 -0.657 -0.659 -0.011 -0.003
Class 5 0.851 0.786 0.934 0.924 -0.718 -0.672 -0.300 -0.297 -0.718 -0.672 -0.267 -0.266
Appendix J Class Wise Results
Here we provide the complete results for the datasets shown in Table 3. In Table 14 we provide results on IMDB dataset on all selected models. In Table 15 we provide results on SST2 dataset on all selected models. In Table 16 we provide results on Emotions dataset on all selected models. In Table 17 we provide results on DBPedia-14 dataset on all selected models.
Table 13: Evaluation of selected models on the AG-News dataset using neuron and range masking techniques. Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance. For GPT-2 50% and for Llama-3.2-3B 30% neurons selected.
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
BERT Class 0 0.945 0.936 0.949 0.927 -0.205 -0.587 0.004 -0.076 -0.198 -0.589 0.007 -0.010
Class 1 0.993 0.988 0.933 0.910 -0.225 -0.659 0.004 -0.077 -0.194 -0.650 0.003 -0.012
Class 2 0.905 0.881 0.962 0.945 -0.300 -0.536 0.014 -0.079 -0.298 -0.542 0.014 -0.009
Class 3 0.949 0.913 0.948 0.935 -0.354 -0.577 0.026 -0.065 -0.353 -0.579 0.025 -0.005
GPT-2 Class 0 0.955 0.951 0.941 0.928 -0.920 -0.926 -0.231 -0.224 -0.919 -0.925 -0.019 -0.008
Class 1 0.986 0.981 0.931 0.917 -0.926 -0.931 -0.253 -0.257 -0.912 -0.916 -0.054 -0.069
Class 2 0.897 0.886 0.960 0.949 -0.696 -0.737 -0.110 -0.132 -0.678 -0.725 -0.097 -0.306
Class 3 0.940 0.916 0.946 0.939 -0.940 -0.916 -0.024 -0.037 -0.887 -0.882 -0.080 -0.510
Llama-3.2-3B Class 0 1.000 0.936 1.000 0.680 -0.995 -0.934 -0.530 -0.427 -0.995 -0.934 -0.345 -0.306
Class 1 1.000 0.742 1.000 0.744 -0.870 -0.680 -0.615 -0.599 -0.875 -0.681 -0.515 -0.503
Class 2 1.000 0.655 1.000 0.773 -0.895 -0.646 -0.795 -0.634 -0.895 -0.646 -0.655 -0.549
Class 3 1.000 0.642 1.000 0.778 -0.975 -0.641 -0.698 -0.630 -0.975 -0.640 -0.420 -0.459
Table 14: Evaluation of selected models on the IMDB dataset using neuron and range masking techniques. Here, Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance.
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
BERT Class 0 0.930 0.908 0.926 0.901 -0.169 -0.352 0.061 -0.066 -0.163 -0.359 0.059 0.035
Class 1 0.926 0.901 0.930 0.908 -0.211 -0.355 0.057 -0.091 -0.206 -0.361 0.056 0.025
GPT-2 Class 0 0.965 0.941 0.940 0.922 -0.935 -0.922 0.050 0.057 -0.905 -0.901 0.055 0.046
Class 1 0.940 0.922 0.965 0.941 -0.620 -0.667 0.005 0.018 -0.610 -0.657 0.015 0.027
Llama-3.2-3B Class 0 1.000 0.619 1.000 0.500 -0.643 -0.448 -0.515 -0.287 -0.640 -0.446 -0.502 -0.278
Class 1 1.000 0.500 1.000 0.619 -0.877 -0.410 -0.273 -0.304 -0.873 -0.409 -0.265 -0.303
Table 15: Evaluation of selected models on the SST2 dataset using neuron and range masking techniques. Here, Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance.
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
BERT Class 0 0.890 0.882 0.930 0.925 -0.058 -0.308 0.029 -0.047 -0.075 -0.329 0.031 0.036
Class 1 0.930 0.925 0.890 0.882 -0.043 -0.318 0.033 -0.045 -0.045 -0.330 0.030 0.050
GPT-2 Class 0 0.950 0.937 0.981 0.978 -0.142 -0.158 0.010 0.012 -0.142 -0.167 0.009 0.010
Class 1 0.981 0.978 0.950 0.937 -0.187 -0.223 0.041 0.053 -0.176 -0.216 0.041 0.046
Llama-3.2-3B Class 0 1.000 0.620 1.000 0.690 -0.532 -0.459 -0.420 -0.424 -0.532 -0.456 -0.404 -0.415
Class 1 1.000 0.690 1.000 0.620 -0.289 -0.379 -0.326 -0.315 -0.284 -0.376 -0.306 -0.301
Table 16: Evaluation of selected models on the Emotions dataset using neuron and range masking techniques. Here, Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance.
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
BERT Class 0 0.960 0.935 0.901 0.851 -0.241 -0.718 0.013 -0.266 -0.222 -0.718 0.012 -0.055
Class 1 0.942 0.904 0.905 0.861 -0.223 -0.691 0.028 -0.254 -0.213 -0.692 0.032 -0.064
Class 2 0.824 0.723 0.926 0.889 -0.371 -0.533 0.016 -0.284 -0.352 -0.534 0.018 -0.115
Class 3 0.927 0.873 0.916 0.876 -0.247 -0.664 0.010 -0.256 -0.240 -0.667 0.012 -0.057
Class 4 0.884 0.837 0.922 0.880 -0.406 -0.646 0.012 -0.251 -0.402 -0.648 0.012 -0.066
Class 5 0.591 0.566 0.929 0.886 -0.303 -0.392 0.004 -0.299 -0.303 -0.397 0.005 -0.090
GPT-2 Class 0 0.969 0.956 0.913 0.903 -0.695 -0.751 -0.125 -0.124 -0.698 -0.749 -0.009 -0.009
Class 1 0.939 0.938 0.925 0.908 -0.879 -0.882 -0.019 -0.009 -0.879 -0.880 -0.016 -0.015
Class 2 0.902 0.872 0.932 0.923 -0.776 -0.736 -0.029 -0.032 -0.780 -0.739 -0.023 -0.028
Class 3 0.910 0.905 0.932 0.921 -0.713 -0.714 -0.006 -0.007 -0.715 -0.716 -0.002 -0.001
Class 4 0.869 0.854 0.938 0.927 -0.754 -0.753 -0.240 -0.248 -0.754 -0.753 -0.127 -0.133
Class 5 0.857 0.798 0.932 0.923 -0.587 -0.601 -0.301 -0.308 -0.587 -0.601 -0.280 -0.289
Llama-3.2-3B Class 0 0.950 0.550 0.782 0.455 -0.950 -0.547 -0.655 -0.408 -0.945 -0.547 -0.571 -0.378
Class 1 0.905 0.498 0.804 0.473 -0.855 -0.495 -0.743 -0.433 -0.867 -0.494 -0.607 -0.404
Class 2 0.785 0.421 0.827 0.483 -0.785 -0.420 -0.771 -0.454 -0.785 -0.420 -0.658 -0.436
Class 3 0.790 0.482 0.833 0.476 -0.760 -0.477 -0.635 -0.423 -0.755 -0.476 -0.544 -0.402
Class 4 0.780 0.487 0.829 0.476 -0.780 -0.486 -0.534 -0.365 -0.780 -0.486 -0.444 -0.324
Class 5 0.536 0.296 0.855 0.498 -0.417 -0.284 -0.751 -0.465 -0.429 -0.282 -0.653 -0.434
Table 17: Evaluation of selected models on the DBPedia-14 dataset using neuron and range masking techniques. Here, Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance.
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
BERT Class 0 0.972 0.966 0.992 0.991 -0.082 -0.702 0.001 -0.014 -0.076 -0.698 0.001 -0.000
Class 1 0.987 0.986 0.991 0.990 -0.030 -0.778 $0$0.017 -0.018 -0.770 $0$0.000
Class 2 0.987 0.985 0.991 0.990 -0.239 -0.814 0.001 -0.018 -0.217 -0.806 0.001 -0.000
Class 3 0.997 0.997 0.990 0.989 -0.008 -0.766 $0$0.019 -0.001 -0.731 $0$0.000
Class 4 0.984 0.983 0.991 0.990 -0.058 -0.777 0.001 -0.018 -0.032 -0.761 $0$0.000
Class 5 0.995 0.995 0.990 0.989 -0.007 -0.795 $0$0.017 -0.001 -0.771 $0$0.000
Class 6 0.975 0.974 0.992 0.991 -0.121 -0.807 $0$0.015 -0.112 -0.803 $0$0.001
Class 7 0.994 0.994 0.990 0.989 -0.028 -0.789 $0$0.017 -0.010 -0.767 $0$0.000
Class 8 1.000 1.000 0.990 0.989 -0.001 -0.808 $0$0.022 $0$0.772 $0$0.000
Class 9 0.999 0.998 0.990 0.989 -0.004 -0.837 $0$0.019 -0.001 -0.811 $0$0.000
Class 10 0.994 0.993 0.990 0.989 -0.025 -0.846 $0$0.016 -0.005 -0.831 $0$0.000
Class 11 0.997 0.997 0.990 0.989 -0.013 -0.751 $0$0.017 -0.001 -0.726 $0$0.000
Class 12 0.990 0.990 0.990 0.989 -0.018 -0.772 $0$0.017 -0.005 -0.755 $0$0.000
Class 13 0.994 0.994 0.990 0.989 -0.009 -0.740 0.001 -0.017 -0.001 -0.721 $0$0.000
GPT-2 Class 0 0.985 0.977 0.990 0.989 -0.860 -0.877 -0.133 -0.136 -0.850 -0.869 -0.002 -0.017
Class 1 0.995 0.992 0.990 0.988 -0.500 -0.567 -0.180 -0.192 -0.460 -0.544 -0.023 -0.024
Class 2 0.985 0.980 0.990 0.989 -0.890 -0.904 -0.189 -0.213 -0.880 -0.902 -0.004 -0.010
Class 3 0.995 0.995 0.990 0.987 -0.900 -0.933 -0.145 -0.143 -0.900 -0.927 -0.008 -0.017
Class 4 0.970 0.969 0.992 0.989 -0.715 -0.773 -0.224 -0.260 -0.695 -0.750 -0.042 -0.062
Class 5 0.995 0.993 0.990 0.988 -0.315 -0.446 -0.127 -0.192 -0.290 -0.432 -0.013 -0.025
Class 6 0.965 0.964 0.992 0.990 -0.925 -0.932 -0.052 -0.062 -0.910 -0.928 -0.006 -0.007
Class 7 1.000 0.998 0.989 0.987 -0.815 -0.865 -0.003 -0.008 -0.775 -0.846 -0.026 -0.057
Class 8 1.000 1.000 0.989 0.987 -0.995 -0.990 -0.148 -0.188 -0.900 -0.932 -0.026 -0.055
Class 9 1.000 1.000 0.989 0.987 -0.975 -0.979 -0.250 -0.268 -0.955 -0.958 -0.020 -0.049
Class 10 0.995 0.993 0.990 0.988 -0.595 -0.685 -0.045 -0.053 -0.590 -0.675 -0.005 -0.011
Class 11 0.985 0.984 0.990 0.988 -0.210 -0.453 -0.094 -0.118 -0.135 -0.396 -0.015 -0.034
Class 12 0.990 0.988 0.990 0.988 -0.930 -0.938 -0.293 -0.309 -0.855 -0.880 -0.013 -0.029
Class 13 1.000 0.999 0.989 0.987 -0.985 -0.986 -0.393 -0.416 -0.945 -0.981 -0.018 -0.044
Llama-3.2-3B Class 0 1.000 0.586 1.000 0.559 -0.990 -0.584 -0.949 -0.473 -0.990 -0.584 -0.823 -0.441
Class 1 1.000 0.533 1.000 0.563 -1.000 -0.528 -0.870 -0.446 -0.970 -0.528 -0.706 -0.371
Class 2 1.000 0.467 1.000 0.568 -0.995 -0.462 -0.963 -0.477 -0.995 -0.461 -0.838 -0.432
Class 3 1.000 0.460 1.000 0.569 -0.995 -0.459 -0.981 -0.486 -0.995 -0.459 -0.815 -0.420
Class 4 1.000 0.828 1.000 0.539 -0.965 -0.809 -0.981 -0.454 -0.955 -0.808 -0.852 -0.412
Class 5 1.000 0.349 1.000 0.568 -1.000 -0.348 -0.882 -0.429 -0.989 -0.347 -0.585 -0.346
Class 6 1.000 0.809 1.000 0.541 -1.000 -0.787 -0.972 -0.449 -1.000 -0.787 -0.736 -0.366
Class 7 1.000 0.599 1.000 0.558 -0.855 -0.588 -0.918 -0.410 -0.860 -0.586 -0.489 -0.274
Class 8 1.000 0.420 1.000 0.572 -1.000 -0.420 -0.957 -0.467 -1.000 -0.420 -0.660 -0.335
Class 9 1.000 0.527 1.000 0.563 -1.000 -0.524 -0.842 -0.435 -0.995 -0.523 -0.552 -0.320
Class 10 1.000 0.505 1.000 0.565 -0.995 -0.503 -0.907 -0.464 -1.000 -0.503 -0.589 -0.322
Class 11 1.000 0.505 1.000 0.565 -0.975 -0.501 -0.862 -0.416 -0.970 -0.501 -0.579 -0.313
Class 12 1.000 0.560 1.000 0.561 -0.980 -0.545 -0.812 -0.417 -0.975 -0.544 -0.496 -0.310
Class 13 1.000 0.587 1.000 0.559 -0.990 -0.584 -0.722 -0.406 -0.985 -0.584 -0.588 -0.337